### v1

# DAWN: Domain-Adaptive Weakly Supervised Nuclei Segmentation via Cross-Task Interactions

Abstract
Weakly supervised segmentation methods have gained significant attention due to their ability to reduce the reliance on costly pixel-level annotations during model training. However, the current weakly supervised nuclei segmentation approaches typically follow a two-stage pseudo-label generation and network training process. The performance of the nuclei segmentation heavily relies on the quality of the generated pseudo-labels, thereby limiting its effectiveness. This paper introduces a novel domain-adaptive weakly supervised nuclei segmentation framework using cross-task interaction strategies to overcome the challenge of pseudo-label generation. Specifically, we utilize weakly annotated data to train an auxiliary detection task, which assists the domain adaptation of the segmentation network. To enhance the efficiency of domain adaptation, we design a consistent feature constraint module integrating prior knowledge from the source domain. Furthermore, we develop pseudo-label optimization and interactive training methods to improve the domain transfer capability. To validate the effectiveness of our proposed method, we conduct extensive comparative and ablation experiments on six datasets. The results demonstrate the superiority of our approach over existing weakly supervised approaches. Remarkably, our method achieves comparable or even better performance than fully supervised methods. Our code will be released in https://github.com/zhangye-zoe/DAWN.
Index Terms:
Nuclei instance segmentation, Weakly supervised learning, Domain adaptation.
I Introduction
Nuclei segmentation [1, 2]plays a vital role in computational pathology, serving as a fundamental step for various nuclei-related tasks. Accurate segmentation enables advanced analysis, such as nuclei classification [3, 4, 5], nuclei counting [6, 7] and nuclei tracing [8, 9, 10]. However, developing supervised nuclei segmentation models in whole slide images (WSIs) is challenging due to the extensive requirement of precise pixel annotation in these images, which consist of gigapixels. At the same time, manual labeling of nucleus boundaries is time-consuming and demands domain expertise. To alleviate the manual burden, a weakly supervised segmentation algorithm is a promising solution, which leverages the inherent image structure and context to achieve segmentation, facilitating more efficient and accurate nuclei analysis.
Figure 1: The weakly supervised nuclei segmentation flows based on point annotation. (a) represents a traditional two-stage segmentation paradigm. The point annotation is employed to generate the pseudo label, which is then used to train the segmentation network. (b) represents our domain-adaptive weakly supervised segmentation method. The point annotation is utilized to train a detection network, which can assist target domain transfer.
Weakly supervised point annotation methods have recently been widely adopted for nuclei segmentation [11, 12, 13], which can decrease manual consumption by a large margin. However, due to inadequate supervision information, these methods generally perform a pre-processing step to generate masks for model training, as shown in Fig. 1 (a). Subsequently, these generated pseudo-labels are utilized to guide the training of the segmentation network. However, this two-stage training approach still faces two significant challenges. Firstly, the pseudo-labels generated through pre-processing often exhibit a coarse granularity and lack precise representation of nuclear morphology. For example, the Voronoi method [11] utilizes distance measurements to differentiate overlapping nuclei, which results in the disruption of boundary morphology, as depicted in Fig. 2 (b). The level set method (LSM) [14] employs color features to simulate nuclear contours. However, it fails to distinguish tissues with similar colors, as illustrated in Fig. 2 (c). Secondly, these two-stage methods still have problems handling the suboptimal pseudo-labels. Although some single-stage optimization methods have been proposed to enhance segmentation performance, such as utilizing boundary mining techniques [15] to improve the quality of pseudo-labels and employing conditional random fields (CRF) [16] for post-processing to rectify segmentation results, these localized improvement methods fail to adapt to the entire training process, resulting in segmentation outcomes that remain unsatisfactory. Therefore, adopting a new training strategy is necessary to mitigate the impact of pseudo-label generation on nuclei segmentation results.
Figure 2: The problems of generated pseudo labels. (a) represents the input image; (b) represents the generated mask using Voronoi [11]; (c) represents the generated mask using LSM [14]; (d) is the ground-truth label of input.
Recent research has demonstrated the potential of domain adaptation in nuclei instance segmentation [17, 18, 19]. They leverage prior information from the source domain and excavate domain-invariant features from the source data to achieve cross-domain task transfer, resulting in remarkable performance. Therefore, it is reasonable to believe that domain-adaptive methods can enhance instance segmentation performance in weakly supervised scenarios. On the one hand, domain adaptation methods do not explicitly generate pseudo-labels for supervising network training [20, 21]. Instead, they enhance the network’s performance in the target domain through domain transfer strategies. It can mitigate the impact of inaccurate pseudo-labels during model training. On the other hand, compared to unsupervised domain adaptation methods, weakly supervised domain adaptation has evident superiority for providing additional supervision information. In the context of weak supervision, incorporating an auxiliary task, such as object detection, can leverage the weak annotations to effectively enhance the segmentation network’s ability to achieve domain transfer [22, 23, 24].
Motivated by these observations, this paper proposes a Domain-Adaptive Weakly supervised Nuclei segmentation method (DAWN). By utilizing point annotation data, DAWN incorporates a cross-task interaction strategy at both the feature and output levels, aiming to enhance the model’s domain transfer capability for accurate nuclei segmentation in weak supervision scenarios. Firstly, we propose a consistent feature constraint (CFC) module, which leverages the task correlation between the detection and segmentation networks to preserve prior knowledge of the target data. Secondly, we introduce a task-level supervision information optimization strategy to address the challenge of the domain mismatch problem. By leveraging the strong localization ability of the detection network, we compensate for the output shortcomings of the segmentation network. Lastly, we devise an interactive training method to achieve progressive transfer of the segmentation network. In detail, we establish mutual supervision between the segmentation and detection tasks. This approach enhances the performance of segmentation predictions by leveraging the complementary aspects of both tasks. Extensive experiment results demonstrate the superiority of the proposed framework. Overall, our contributions can be summarized as follows:
  • We propose a novel domain-adaptive weakly supervised nuclei segmentation framework that departs from traditional two-stage weakly supervised methods. Our approach leverages a detection task with point annotations and domain transfer strategies to address the issue of inaccurate pseudo-label generation.
  • We devise a consistent feature constraint module, which utilizes prior information of the source domain to improve the nuclei segmentation performance of target domain.
  • We develop a combined pseudo-label optimization method incorporating the output information from the segmentation and the detection networks. Besides, through an iterative training strategy, this module enhances the domain transfer capability at the task level.
  • We propose an interactive supervised training method that leverages optimized pseudo-labels to simultaneously supervise the training of segmentation and detection networks. This approach facilitates the learning of complementary features between the two networks.
  • We validate our method on six datasets. Extensive experiments indicate that our method outperforms existing weakly supervised methods and achieves similar or even better performance with fully supervised methods.
II Related Work
II-A Fully Supervised Nuclei Segmentation
Nuclei segmentation is an essential workflow, and analysis based on nuclei can provide a basis for downstream tasks of computational pathology. In the past, watersheds [25] or Otsu threshold[26] usually led to over-segmentation or under-segmentation problems. With the successes of U-Net[8] in the cell tracing task, the variations of U-Net are emerging [27, 28, 29] and widely used in the nuclei segmentation. At present, nuclei segmentation methods mainly focus on solving the nuclei overlapping and unclear staining problems. To solve nuclei overlapping, HoverNet[30] calculates the horizontal and vertical gradients to supervise training. Cellpose[31] uses simulated diffusion and spatial gradients to trace instances to which each pixel belongs. CDNet[32] employs the direction difference map to distinguish different instances of overlapping regions. Besides, DCAN [33] adds the contour category to the semantic prediction task to separate instances. To solve the unclear staining problem, some staining separation methods [34, 35] are proposed to extract nuclei staining components. In addition, Sams-net [36] and Triple-Unet [37] use H-channel of staining separation images to enhance the nuclei feature representation ability. Although these methods achieve good progress in nuclei segmentation, they need a large amount of pixel-level manual annotation, which is expensive and labor-intensive. On the contrary, weakly supervised methods provide an effective solution to reduce the reliance of model development on annotations. They allow training models with coarse annotations, reducing time and manual costs.
II-B Weakly Supervised Segmentation
Existing weakly supervised segmentation labels mainly include image-level, bounding-box, and point annotations. Among that, the models using image-level annotations typically rely on the Class Activation Map (CAM) [38] to generate pseudo labels. PRM[39] utilizes peak backpropagation from CAMs to generate highly informative object regions. SC-CAM[40] accelerates the integrity of CAMs using the unsupervised clustering sub-category information. RRM[41] uses the CAM and reliable score evaluation to generate finer pseudo labels. The methods utilizing bounding-box annotations mainly focus on the saliency models. BoxCaseg[42] employs the salient segmentation head to enhance the object localization ability. BoxTeacher[43] uses noise-aware pixel loss and noise-reduced affinity loss to optimize the pseudo labels. Due to the small nuclei morphology, point labels are widely adopted as the supervision information in pathological images. WeakSeg [16] develops Voronoi splitting and clustering to generate pseudo labels. After that, Voronoi is widely used to generate segmentation masks[11, 44, 45, 13]. Besides, [46] establishes a propagating detection map to segment fluorescence images. [12] adopts a co-training strategy to segment phase images. Even though these weakly supervised nuclei segmentation methods have achieved good results, they often lead to poor shape prediction due to the need for more morphological consideration. In this paper, we improve the nuclear morphology through a pre-training model and domain adaptation strategies.
II-C Domain Adaptation in Medical Image Analysis
Pre-training methods [47] can bring convenience to small-sample task training. However, directly using pre-trained models often leads to poor performance due to the distribution differences of samples. Therefore, some domain adaptation methods are proposed to address the problem of domain transfer. [48] uses the image translation method to align the pixel value distribution, reducing the gap between the source domain and the target domain. DIGA[49] develops a distribution adaptation and a semantic adaptation module to enable model adaptation jointly. Inspired by natural scenario methods[50], domain adaptation is also used in medical image analysis[17]. DARCNN[19] implements domain adaptation from natural scene images to hematoxylin-eosin (H&E) images, exploiting domain-invariant and domain-specific features. CyC-PDAM[51] transfers the model from the fluorescence images domain to the H&E images domain. CAPL-Net [52] implements domain adaptation in H&E images from different sources. SAMDA [53] combines the Segment Anything Model (SAM) with nnUNet [54] in the embedding space to achieve a few-shot domain adaptation. However, existing models mainly utilize unsupervised methods for domain adaptation, often leading to poor performance due to the lack of understanding of target domain features. In this paper, we leverage auxiliary tasks to establish connections between the source domain model and the detection task, which promotes the transfer efficiency of the source model to the target domain.
Figure 3: The schematics of our proposed DAWN framework. (a) represents the pre-trained segmentation network in source domain; (b) expresses the generation of extended point annotation using Gaussian distribution; (c) and (d) represent target domain transfer process.
III Methods
III-A Framework Overview
Detection and segmentation are closely related tasks that share similar semantic information, such as the location and texture of the object, which can be utilized to assist in instance detection or segmentation. In multiple-task learning scenarios, detection and segmentation networks are usually combined to improve the performance of related tasks [55, 56]. At that point, under the domain adaptation background, we design a detection task based on point annotations to assist the segmentation network domain transfer. By enhancing the adaptability of the segmentation network to target domain data, the nuclei segmentation accuracy can be improved in weakly supervised scenarios.
This paper establishes cross-task interaction strategies at both the feature and task levels. The feature interaction is designed to preserve the prior segmentation information related to the target domain. At the same time, the task interaction is used to compensate for the shortcomings of the segmentation output through combined pseudo-label optimization and interactive supervision methods. Overall, the flowchart of our proposed DAWN is depicted in Fig. 3.
We firstly pre-train a segmentation network in source domain as shown in Fig. 3 (a). To overcome the domain shift problem caused by data distribution variations, we design a consistent feature constraint (CFC) module and a combined pseudo-label optimization (CPL) module as shown in Fig. 3 (c)-(d). In addition, to promote domain transfer ability and efficiency, we employ an interactive supervision training (IST) method. According to these constraints, our training loss function is formulated as follows:
$Loss=L_{det}+\alpha L_{fea}+\beta L_{dyn},$ (1)
where $L_{det}$ is the loss of detection network, $L_{fea}$ is the consistent feature constraint loss, and $L_{dyn}$ is the dynamic supervision loss. $\alpha$ and $\beta$ are used to balance the weight of each loss. It should be noted that $L_{det}$ corresponds to a fully-supervised learning task.
III-B Consistent Feature Constraint
The semantic features encoded by detection and segmentation networks are correlated and complementary, which provides a basis for adaptive learning of the segmentation network with the help of an auxiliary detection network. The source domain dataset is defined as $D_{S}=\{(x_{s},y_{s}^{f})\}_{s=1}^{m}$ and the target domain dataset is defined as $D_{T}=\{(x_{t},y_{t}^{p})\}_{t=1}^{n}$. $x_{s}$ and $x_{t}$ represent the input image of the source and target domain, respectively. $y_{s}^{f}$ represents fine pixel-level label and $y_{t}^{p}$ represents point label. $m$ and $n$ represent the number of images in the source and target domain, respectively.
On source domain data $D_{S}$, we firstly pre-train a segmentation network as shown in Fig. 3 (a) and Stage 1 of Algorithm 1. We adopt HoverNet [30] as a segmentation network in this step, and its supervision information consists of three components: foreground probability $p_{s}$, horizontal distance $h_{s,x}$, and vertical distance $h_{s,y}$. Using the pixel-level annotation $y_{s}^{f}$, the supervision images can be obtained through pre-processing, as shown in Fig. 3 (a). In this step, the loss function is formulated as follows:
$\begin{split}L_{pre}&=CE(\hat{p}_{s},p_{s})+Dice(\hat{p}_{s},p_{s})\\ &+MSE(\hat{h}_{s,x},h_{s,x})+MSE(\hat{h}_{s,y},h_{s,y}),\end{split}$ (2)
where $\hat{p}_{s}$, $\hat{h}_{s,x}$, and $\hat{h}_{s,y}$ represent the foreground probability prediction, horizontal distance prediction, and vertical distance prediction. CE, Dice, and MSE represent cross-entropy loss, dice loss, and mean square error loss.
In the following domain adaptation process, we jointly train the segmentation and detection networks, and the process is shown in Fig. 3 (c) and Stage 2 of Algorithm 1. Firstly, the well-trained segmentation network is utilized to initialize the target domain segmentation model. Then, we map the output feature $f_{seg}$ of encoder into $f_{seg}^{c}$ by a convolution layer and a fully connection layer:
$f_{seg}^{c}=FC(Conv(f_{seg})),$ (3)
where $FC$ and $Conv$ represent fully connected and convolution layers, respectively. Finally, the output feature $f_{seg}^{c}$ is used to compute the consistent feature constraint. It should be noted that the detection network is trained in the target domain, and the segmentation network is initialized in the source domain. Hence, the CFC can force the segmentation network to gradually close to the target domain.
Considering single central pixel of nuclei just represents nuclei position while ignoring nuclei texture, it is hard to train an effective detection network using point label $y_{p}^{t}$ [57, 58]. Thus, we employ extended Gaussian distribution [16] to model nuclei morphology. The extended Gaussian encoding process is shown in Fig. 3 (b) and can be formulated as follows:
$\begin{split}M_{g,i}=\left\{\begin{array}[]{ll}exp(-\frac{{D_{i}}^{2}}{2{\sigma}^{2}}),&\;D_{i}\leq r_{1},\\ 0,&\;r_{1}<D_{i}\leq r_{2},\\ -1,&otherwise,\end{array}\right.\end{split}$ (4)
where $D_{i}$ represents the distance between pixel $i$ and nearest point annotation, $r_{i}$ represents initialized nuclei radius, the region between $r_{1}$ and $r_{2}$ is background, $\sigma$ is the discrete degree of Gaussian distribution. In $M_{g,i}$, the pixel corresponding to $-1$ does not participate in the training. Employing the generated encoding, a supervised detection model is trained as follows:
$L_{det}=\frac{1}{|\Omega|}\sum_{i\in\Omega}w_{i}(\hat{q}_{i}-M_{g,i})^{2},$ (5)
where $\Omega$ represents the pixel set participating training, i.e. $\Omega=\{i|0<D_{i}\leq r_{2}\}$; $w_{i}$ is the training weight of pixel $i$. In order to pay more attention to the foreground, we set $w_{i}$ equals $10$ when $D_{i}<r_{1}$; otherwise, $w_{i}$ equals $1$. The extended Gaussian encoding not only encodes the tissue texture but can also learn nuclei shape gradually, assisting segmentation network optimization.
Similar to the above segmentation network, the middle layer output $f_{det}$ of the detection network is mapped into $f_{det}^{c}$ to calculate CFC as shown in Fig. 3 (c) and the feature constraint loss $L_{fea}$ is calculated as follows:
$L_{fea}=\frac{1}{|dim|}{\|f_{seg}^{c}-f_{det}^{c}\|}^{2}_{2},$ (6)
where $dim$ represents the dimension of feature encoding. $f_{seg}^{c}$ and $f_{det}^{c}$ are mapped features of segmentation and detection networks. In this process, the CFC module forces the two networks to excavate the essential task correlation features, enhancing domain adaptation efficiency.
Algorithm 1 The Whole Training Process of Our DAWN
1:  Stage 1. Pre-train segmentation network
2:  Input: Source domain data $D_{s}$; number of epoch $e_{s}$;
3:  Output: Network weight $W_{s}$;
4:  for epoch = 1,…, $e_{s}$ do
5:     Calculate loss $L_{pre}$ in Eq. 2 and update network;
6:     Output: $\hat{p}_{s}$, $\hat{h}_{s,x}$, $\hat{h}_{s,y}$;
7:  end for
8:  Save weights: $W_{s}$.
9:  Stage 2. Joint training networks
10:  Input: Target domain data $D_{t}$; Point annotation $y_{t}^{p}$; initialized weight $W_{s}$;
11:  Initialize: $r=4$; $e_{t}=20$; Initialize segmentation network with $W_{s}$;
12:  Output: Instance segmentation results $\hat{y}_{t}$;
13:  for iter=1, …, $r$ do
14:     for epoch=1,…,$e_{t}$ do
15:        Calculate loss: $L_{det}+\alpha L_{fea}+\beta L_{dyn}$ in Eq. 1 and update network;
16:     end for
17:     Prediction outputs: $\hat{p}_{t}^{seg}$, $\hat{q}_{t}^{det}$;
18:     Update pseudo-label: $M_{pse}$leftarrow CPL($\hat{p}_{t}^{seg}$, $\hat{q}_{t}^{det}$, $y_{t}^{p}$);
19:     Training networks with $M_{pse}$ in Eq. 9;
20:     Post-process: $\hat{y}_{t}$leftarrow Post($\hat{p}_{t}$, $\hat{h}_{t,x}$, $\hat{h}_{t,y})$;
21:  end for
III-C Combined Pseudo-label Optimization
Due to domain shift resulting from nuclei morphology and size variations across different organs, there may be an under-segmentation problem in the target domain prediction $\hat{y}_{t}^{seg}$. To solve the problem, we devise a combined pseudo-label (CPL) optimization method by combining the outputs of the segmentation and detection networks. This method utilizes the superior object localization capabilities of the detection network and effectively compensates for the flaw of the segmentation network. The schematic diagram of the method is depicted in Fig. 4.
The CPL method comprises two key components: mask fusion and distance filter. In mask fusion, the prediction outputs from both networks are combined to achieve comprehensive nuclei segmentation. The distance filter is then applied to refine the fused results by eliminating incorrectly segmented nuclei. The resulting outputs from the CPL serve as the ground truth annotations for supervising the subsequent training.
III-C1 Mask Fusion
Based on the prediction output $\hat{q}_{t}^{det}$ of the detection network, we employ a threshold filter to preserve the high confidence region of prediction $\hat{q}_{t}^{det}$ as shown in Fig. 4 (a) and obtain the initialized detection $M_{det}$ as shown in Fig. 4 (b), which represents the nuclei region:
$M_{det}=I_{[\hat{q}_{t}^{det}>\theta]}\in\{0,1\}^{N\times N},$ (7)
where $\hat{q}_{t}^{det}$ represents the probability output of detection model, $\theta$ is the threshold value, $M_{det}$ represents generated nuclei region. At the same time, we transform the probability output $\hat{p}_{t}^{seg}$ of the segmentation network, obtaining the initialized $M_{ini}$, which has the under-segmentation problem as shown in Fig. 4 (c). Then, we calculate a union set for $M_{ini}$ and $M_{det}$ as shown in Fig. 4 (d):
$M_{uni}=M_{ini}\cup M_{det}.$ (8)
In mask fusion, by utilizing the strong identification ability of $M_{det}$, we solve the under-segmentation problem led by the domain shift. The generated $M_{uni}$ contains the prediction result of $M_{ini}$ and a new mask in $M_{det}$. It is worth noting that although detection output $M_{det}$ has an irregular nuclei shape, it will be optimized in the interactive supervision process, which will be stated in the following subsection.
Figure 4: The schematic diagram of combined pseudo-label optimization. (a) is probability output of detection network; (b) is detection mask through filtering; (c) is segmentation mask; (d) is combined map of detection and segmentation outputs; (e) is point annotation; (f) is generated pseudo-label.
III-C2 Distance Filter
We propose a distance filter method to solve incorrect segmentation in $M_{uni}$. This method utilizes point labels to locate the preserved nuclei and employs a distance-based metric to eliminate incorrectly segmented nuclei. Specifically, our method is described as follows.
Firstly, we use the point annotation $y_{t}^{p}$ of the target domain data to locate the nuclei in the fused image $M_{uni}$. As illustrated in Fig. 4 (f), the red dots represent the annotated points. Next, we calculate the distance from each pixel to its nearest annotation point and choose pixels within a certain distance. In this process, the distance is set as $d$, and the pixels that have a distance smaller than $d$ are considered foreground, while other regions are treated as background. After the distance filter, we obtain the updated pseudo-label $M_{pse}$ as shown in Fig. 4 (g). The pseudo label $M_{pse}$ is a binary map; in the map, “1” represents the foreground, and “0” represents the background, which is used to guide the interactive training.
III-D Interactive Supervision Training Method
This sub-section proposes a dynamic pseudo-labels optimization method to facilitate domain adaptation in the segmentation network. Specifically, the pseudo-labels generated from the CPL module in the $(i-1)^{th}$ round will be viewed as supervision information for the training of segmentation and detection networks in the $i^{th}$ round. By continuously optimizing the segmentation network, this method helps the progressive domain transfer of the network. Our interactive supervision process is shown in Fig. 3 (c)-(d) and Algorithm 1 Line 13-16.
In the $(t-1)^{th}$ round, the CPL module outputs the pseudo-labels, denoted as $M_{pse}$. The pseudo-label $M_{pse}$ combines the previous outputs of the segmentation and detection networks. Therefore, $M_{pse}$ is viewed as the supervision information for training the new round of the segmentation network, effectively addressing the under-segmentation issue in the segmentation results. At the same time, point annotation-based methods are used to supervise the detection network, which does not fully consider the morphological information of nuclei. Hence, we employ pseudo-labels to supervise the training of the detection network, which can accelerate the detection network to learn the nuclear morphology. During the training process, $M_{pse}$ always contains the prediction information from both the segmentation and detection networks. Thus, this method can be considered as an interactive supervision process, and the loss function for this method is shown below:
$L_{dyn}=CE(M_{pse},\hat{p}_{t}^{seg})+MSE(M_{pse},\hat{q}_{t}^{det}),$ (9)
where $M_{pse}$ is the pseudo-label generated by the CPL module, $\hat{p}_{t}^{seg}$ and $\hat{q}_{t}^{det}$ represent the predicted output of segmentation and detection networks in the target domain.
By using the pseudo-labels to supervise the network training in a dynamic and progressive manner, our approach promotes the effectiveness of domain adaptation and enhances the performance of the segmentation network.
TABLE I: Hyper-parameter settings.
Dataset $r_{1}$ $r_{2}$ $\sigma$ $\theta$ d
TNBC 11 22 2.75 0.2 25
CryoNuSeg 11 22 2.75 0.2 25
Lizard 8 16 2 0.2 25
ConSeP 11 22 5 0.5 25
IV Experiments
IV-A Datasets
In this subsection, we briefly introduce the PanNuke, MoNuSeg used in the source domain (♣), and TNBC, CryoNuSeg, ConSeP used in the target domain (♠) in this paper. The PanNuke, MoNuSeg, TNBC, CryoNuSeg, and ConSeP datasets are sampled at ×40 magnification. The lizard dataset is sampled at ×20 magnification.
PanNuke[59] comes from TCGA (The Cancer Genome Atlas), which consists of 205,343 labeled nuclei from 19 different organs in more than 2,000 visual fields.
MoNuSeg[60] comes from TCGA, which consists of 30 histopathology images from 7 different organs.
TNBC[61] comes from Curie Institute, which consists of 50 histopathology images from 11 breast cancer patients.
Lizard[62] comes from UHCW (University Hospital Coventry and Warwickshire) and TCGA, which consists of 291 image regions from colorectal cancer patients.
CryoNuSeg[63] comes from TCGA. The dataset consists of 30 histopathology images from 10 organs.
ConSeP[30] comes from UHCW, which consists of 41 histopathology images from colon cancer patients.
IV-B Implementation Details and Evaluation Metrics
Our DAWN uses a U-Net framework with a ResNet [64] backbone as a detection network. Meantime, a ResNet encoder combined with a DenseNet decoder is used as a segmentation network. Besides, all experiments are run with PyTorch on an Nvidia RTX 3090 GPU. We set loss weights $\alpha=0.1$, $\beta=0.15$ used in Eq. 1. The hyperparameters used Eq. 4 and Eq. 7 are shown in Table I. $r_{1}$ is set according to the mean nuclei radius of the validation set, and $r_{2}$ is set to double $r_{1}$ for balancing foreground and background pixels. $\theta$ is set according to the detection network output on the validation set. We evaluate our experiments over metrics of DICE, AJI[60], DQ[65], SQ[65] and PQ[65]. Among those, AJI, DQ, SQ, and PQ are instance-level metrics. DQ evaluates the detection quality, SQ evaluates segmentation quality, and PQ is the product of the DQ and SQ. Throughout all the tables in this paper, we bold the best and underline the second best.
TABLE II: The comparison of different weakly supervised and fully supervised methods.
  Datasets Methods DICE AJI DQ SQ PQ Datasets Methods DICE AJI DQ SQ PQ
  DDTNet[14] 0.789 0.665 0.807 0.745 0.601 DDTNet[14] 0.602 0.236 0.350 0.619 0.259
BoNuS[15] 0.833 0.712 0.798 0.743 0.635 BoNuS[15] 0.623 0.264 0.364 0.617 0.309
WeakSeg[16] 0.811 0.647 0.804 0.764 0.616 WeakSeg[16] 0.714 0.381 0.403 0.676 0.334
MaskGANet[45] 0.751 0.597 0.744 0.612 0.454 MaskGANet[45] 0.603 0.197 0.256 0.584 0.134
C2FNet[13] 0.754 0.605 0.763 0.682 0.505 C2FNet[13] 0.597 0.213 0.286 0.600 0.163
TNBC PseEdgeNet[66] 0.682 0.497 0.688 0.547 0.394 Lizard PseEdgeNet [66] 0.545 0.140 0.229 0.534 0.105
(a) WeakSAM[67] 0.834 0.697 0.795 0.782 0.620 (c) WeakSAM[67] 0.643 0.279 0.314 0.608 0.267
U-Net[8] 0.783 0.692 0.754 0.750 0.617 U-Net[8] 0.716 0.365 0.436 0.633 0.353
HoverNet[30] 0.815 0.662 0.793 0.811 0.644 HoverNet[30] 0.803 0.465 0.573 0.780 0.453
CDNet[32] 0.847 0.724 0.840 0.819 0.684 CDNet[32] 0.708 0.507 0.630 0.764 0.487
DCAN[33] 0.704 0.497 0.623 0.661 0.433 DCAN[33] 0.716 0.290 0.435 0.632 0.277
Dist[61] 0.836 0.681 0.770 0.792 0.611 Dist[61] 0.753 0.520 0.588 0.676 0.423
Micro-Net[68] 0.813 0.673 0.772 0.795 0.615 Micro-Net[68] 0.776 0.519 0.615 0.759 0.451
DAWN (Ours) 0.856 0.728 0.843 0.820 0.693 DAWN (Ours) 0.728 0.403 0.510 0.736 0.377
  Datasets Methods DICE AJI DQ SQ PQ Datasets Methods DICE AJI DQ SQ PQ
  DDTNet[14] 0.696 0.440 0.618 0.668 0.389 DDTNet[14] 0.598 0.295 0.282 0.452 0.186
BoNuS[15] 0.693 0.431 0.607 0.645 0.399 BoNuS[15] 0.651 0.354 0.378 0.675 0.380
WeakSeg[16] 0.682 0.410 0.579 0.649 0.357 WeakSeg[16] 0.646 0.366 0.363 0.669 0.391
MaskGANet[45] 0.647 0.393 0.595 0.654 0.380 MaskGANet[45] 0.653 0.367 0.401 0.696 0.374
C2FNet[13] 0.643 0.384 0.581 0.573 0.332 C2FNet[13] 0.749 0.464 0.497 0.706 0.398
CryoNuSeg PseEdgeNet[66] 0.620 0.321 0.507 0.529 0.306 ConSeP PseEdgeNet [66] 0.569 0.226 0.245 0.307 0.171
(b) WeakSAM[67] 0.754 0.480 0.627 0.685 0.449 (d) WeakSAM[67] 0.669 0.378 0.384 0.680 0.401
U-Net[8] 0.697 0.469 0.546 0.731 0.403 U-Net[8] 0.761 0.499 0.569 0.725 0.434
HoverNet[30] 0.804 0.526 0.655 0.754 0.495 HoverNet[30] 0.837 0.513 0.636 0.767 0.492
CDNet[32] 0.776 0.539 0.660 0.754 0.499 CDNet[32] 0.835 0.541 0.652 0.771 0.514
DCAN[33] 0.766 0.335 0.379 0.681 0.258 DCAN[33] 0.745 0.312 0.439 0.683 0.300
Micro-Net[68] 0.690 0.435 0.463 0.701 0.327 Micro-Net[68] 0.805 0.516 0.602 0.755 0.457
Dist[61] 0.792 0.518 0.595 0.721 0.432 Dist[61] 0.793 0.512 0.587 0.716 0.422
DAWN (Ours) 0.804 0.508 0.637 0.744 0.476 DAWN (Ours) 0.805 0.509 0.627 0.759 0.477
IV-C Comparison with Weakly Supervised Methods
We compare our proposed domain-adaptive weakly supervised nuclei segmentation method with seven other models on four datasets, and the comparison methods consist of DDTNet[14], BoNuS[15], WeakSeg[16], MaskGANet[45], C2FNet[13], PseEdgeNet[66], and WeakSAM[67]. The experiment results are shown in Table II. In the table, the best weakly supervised method is bold with blue text, and the second best weakly supervised method is underlined with blue text. For fair comparison, for the two-stage methods WeakSeg, BoNuS and DDTNet, we use the same pseudo labels as our DAWN. In detail, the pre-trained segmentation network is test on target dataset and the prediction outputs is used as pseudo labels. In the experiment, we use PanNuke as source domain dataset.
IV-C1 Comparisons on Breast Cancer TNBC
We first validate our proposed domain-adaptive weakly supervised nuclei segmentation on a breast cancer dataset, and the experimental results are shown in Table II. From this table, we observe that our proposed method outperforms existing state-of-the-art weakly supervised segmentation methods. Compared with the best method, our DAWN achieves improvements of 2.2% and 1.6% in terms of DICE and AJI metrics, respectively. Moreover, our method exhibits significant advantages over traditional two-stage weakly supervised models such as MaskGANet and WeakSeg. Our method surpasses MaskGANet by at least 10% and WeakSeg by at least 4.0% on all five metrics.
Furthermore, we conduct visualization comparison experiments as shown in the first two rows of Fig. 5. From the figure, it can be observed that our proposed DAWN, WeakSeg, DDTNet, and BoNuS can effectively segment nuclei. However, methods like MaskGANet and PENet exhibit weaker morphological learning capabilities. We attribute this to pre-trained segmentation networks, where the original domain’s segmentation network contains a wealth of prior morphological information that improves nuclei segmentation performance.
IV-C2 Comparisons on Pan-Cancer CryoNuSeg
We evaluate the performance of our proposed domain adaptation method for weakly supervised segmentation on a challenging pan-cancer dataset called CryoNuSeg. It is worth noting that the CryoNuSeg dataset exhibits significant variations in nuclei morphology, making it a problematic dataset for training. Our experimental results are presented in Table II (b). From these results, it can be observed that our proposed DAWN outperforms the best-performing method, WeakSAM, by 5.0% and 2.8% in terms of DICE and AJI metrics, respectively. Furthermore, compared to the second-best method, DDTNet, our DAWN achieves approximately 10% improvement in DICE, AJI, and PQ metrics.
We perform visual comparison experiments, and the results are shown in the third and fourth rows of Fig. 5. The results of WeakSeg, MaskGANet, C2FNet, and PENet show that these methods exhibit poorer segmentation results when faced with significant morphological variations of nuclei. In contrast, our method still achieves outstanding performance. We attribute this phenomenon to the fact that two-stage methods struggle to adapt their pseudo-labels to the variations in staining and morphology, leading to lower segmentation performance. However, our method utilizes a domain transfer strategy that leverages prior knowledge from the source domain and introduces a progressive domain adaptation method to adapt to the target domain’s image characteristics, resulting in better segmentation results.
IV-C3 Comparisons on Colorectal Cancer Lizard and ConSeP
We further validate the performance of our proposed method on colorectal cancer pathological images. From the Lizard results in Table II (c), we observe that existing weakly supervised methods perform poorly, especially C2FNet, MaskGANet, and PseEdgeNet, with PQ scores of only 0.134, 0.163, and 0.105, respectively. The phenomenon is because the Lizard dataset is sampled at the magnification of 20, resulting in densely distributed and small-sized nuclei, posing challenges for the segmentation task. Nevertheless, our method still achieves improvements compared to the best-performing WeakSeg. We observe improvements of 1.4% and 2.2% for DICE and AJI metrics, respectively. Additionally, from the fifth and sixth rows of Fig. 5, it can be seen that the nuclei in the Lizard dataset are small and exhibit dense distribution, and most models struggle to produce satisfactory results. However, our approach can better identify the nuclei compared to other methods.
From the ConSeP results in Table II (d), it can be observed that our method achieves significant improvements compared to other weakly supervised models. Compared to C2FNet, our DAWN model achieves improvements of 5.6% and 4.5% in terms of DICE and AJI, respectively. From the visual results in the last two rows of Fig. 5, it can be seen that the nuclei in the ConSeP dataset often have blurry morphology and dense distribution. Other methods struggle to identify the boundaries and distinguish individual nuclei instances accurately. In contrast, our method partially addresses this issue and somewhat improves the segmentation performance.
Figure 5: The visualization comparison between weakly supervised methods and our proposed DAWN.
IV-D Comparison with Fully Supervised Methods
To compare the performance difference between our proposed weakly supervised segmentation method and existing fully supervised segmentation methods, we compare our DAWN with six state-of-the-art models, including U-Net [8], HoverNet [30], CDNet [32], DCAN [33], Micro-Net [68], and Dist [61]. In the experiment, we use PanNuke as the source domain and TNBC and ConSeP as the target domains. For a fair comparison, all comparison models are pre-trained in PanNuke and then fine-tuned in the four datasets. The results are shown in Table II. In the table, the best fully supervised method is bold with red text, and the second best fully supervised method is underlined with red text.
IV-D1 Quantitative Comparisons
From Table II, it can be seen that our method achieves comparable performance to fully supervised methods on the TNBC dataset. In the DICE, our method outperforms the best-performing CDNet by 0.4% and significantly surpasses other traditional supervised methods such as U-Net and DCAN. On the CryoNuSeg and ConSeP datasets, although our method does not achieve the best performance, it is only 1.8% and 0.4% lower than the second-best HoverNet method in terms of AJI. However, our method still performs poorly on the Lizard dataset. While it outperforms U-Net and DCAN, there is still a performance gap compared to other methods. We attribute this phenomenon to the limitations of our domain adaptation method when dealing with densely distributed nuclei data, and using point annotations for domain transfer still poses challenges.
IV-D2 Qualitative Comparisons
Furthermore, we provide comparison visualization in Fig. 6. From the figure we can see our weakly supervised segmentation method achieves comparable performance with supervised methods. Surprisingly, our method even outperforms Micro-Net and U-Net. This indicates an accurate pseudo label initialization and an elaborate pseudo label optimization also provide fine supervision information although without pixel-level mask.
Figure 6: The segmentation comparison between fully supervised methods and our proposed DAWN. Blue stands for correctly true positive pixels, red stands for false positive pixels and green stands for false negative pixels.
IV-E Domain Adaptation Evaluation
We conduct a domain adaptation comparison with three domain adaptation methods: SIFA[17], PDAM[18] and SFDA-CellSeg[69]. The comparison experiment is designed from three perspectives.
IV-E1 PanNuke as Source dataset
To compare the different domain adaptation methods, we use a large pan-cancer dataset, PanNuke, as the source domain data and TNBC (breast cancer) and CryoNuSeg (pan-cancer) as the target domain data for the experiment. The experiment results are shown in Table III. From this table, we can observe that our method achieves significant improvements both on TNBC and CryoNuSeg datasets. Regarding the AJI metric, our method outperforms the second-best method by 4.63% and 4.49% on TNBC and CryoNuSeg, respectively.
TABLE III: The domain adaptation comparison experiments from source domain PanNuke to target domain TNBC and CryoNuSeg.
[…]
DICE AJI DQ SQ PQ
source only 0.8216 0.6764 0.8158 0.8094 0.6610
SIFA [17] 0.7934 0.6520 0.7827 0.7493 0.6291
PDAM [18] 0.8192 0.6733 0.8104 0.7831 0.6577
SFDA-CellSeg [69] 0.8340 0.6817 0.8025 0.7933 0.6715
Ours 0.8561 0.7280 0.8425 0.8202 0.6925
[…]
DICE AJI DQ SQ PQ
source only 0.7857 0.4699 0.6121 0.7467 0.4667
SIFA [17] 0.7437 0.4526 0.5731 0.7012 0.4011
PDAM[18] 0.7855 0.4748 0.6030 0.7381 0.4461
SFDA-CellSeg [69] 0.7912 0.4630 0.6121 0.7277 0.4480
Ours 0.8041 0.5079 0.6374 0.7439 0.4760
TABLE IV: The domain adaptation comparison experiments from source domain MoNuSeg to target domain TNBC and CryoNuSeg.
[…]
DICE AJI DQ SQ PQ
source only 0.7574 0.6323 0.7515 0.7333 0.6017
SIFA [17] 0.7530 0.6289 0.7304 0.7157 0.5858
PDAM [18] 0.7873 0.6421 0.7511 0.7543 0.6117
SFDA-CellSeg [69] 0.8007 0.6597 0.7404 0.7484 0.6301
Ours 0.8294 0.6933 0.8074 0.8085 0.6679
[…]
DICE AJI DQ SQ PQ
source only 0.7611 0.4534 0.5698 0.7221 0.4002
SIFA [17] 0.7404 0.4421 0.5671 0.6924 0.3988
PDAM [18] 0.7679 0.4654 0.5899 0.7121 0.4234
SFDA-CellSeg [69] 0.7687 0.4521 0.6004 0.7133 0.4287
Ours 0.7864 0.4845 0.6122 0.7216 0.4498
IV-E2 MoNuSeg as Source dataset
To validate the influences of data scale, we conducted domain adaptation experiments on a smaller pan-cancer dataset, MoNuSeg. The experiment results are shown in Table IV. When using MoNuSeg as the source domain data, the AJI metric of our method surpasses the second-best method by 3.36% and 3.24% on TNBC and CryoNuSeg target data, respectively. This indicates that our method is effective on small-scale datasets. Besides, comparing the results in Table III and Table IV, the latter’s performance significantly decreases. On the CryoNuSeg data, the DICE metric for the three comparison methods decreased by 0.33%, 2.06%, and 2.25%, while our proposed method decreased by only 1.77%, showing a smaller decline compared to the PDAM and SFDA-CellSeg. The phenomenon suggests that our method is less affected by the data scale disturbance.
IV-E3 Cross-cancer Domain Adaptation
We conduct cross-cancer domain adaptation experiments on the ConSeP (colon cancer) and TNBC (breast cancer) datasets, and the experiment results are shown in Table V. We can observe that our method achieves the best performance among these methods. Particularly on the TNBC, the AJI surpasses the second-best method by 4.20%. These experiments demonstrate the significant advantage of our method. Based on this phenomenon, we can conclude that our method is least affected by the source domain data. Our method can still exploit essential prior information for different cancer types to enhance the segmentation performance.
Figure 7: The comparison of before and after domain adaptation. Blue circles indicate instances which are not recognition in source only method; Red and white circles represent accurate instance segmentation in our DAWN; Yellow arrow represents more smoother boundary of our DAWN, which is enlarged in bottom-right corner.
TABLE V: The cross-cancer domain adaptation experiments from source domain ConSeP (TNBC) to target domain TNBC (ConSeP).
[…]
DICE AJI DQ SQ PQ
source only 0.7544 0.6035 0.7144 0.6961 0.5736
SIFA [17] 0.7101 0.5414 0.5528 0.7012 0.5413
PDAM [18] 0.7302 0.5661 0.5757 0.7279 0.5725
SFDA-CellSeg [69] 0.7582 0.6081 0.6033 0.7322 0.5921
Ours 0.8285 0.6634 0.7894 0.7855 0.6453
[…]
DICE AJI DQ SQ PQ
source only 0.6879 0.4391 0.5273 0.6400 0.3924
SIFA [17] 0.6786 0.4076 0.4400 0.6592 0.3657
PDAM [18] 0.6829 0.4324 0.4590 0.6782 0.3712
SFDA-CellSeg [69] 0.6913 0.4217 0.4484 0.6520 0.3817
Ours 0.7743 0.4607 0.5549 0.6981 0.4168
IV-E4 Qualitative Comparisons
We provided visualization comparisons of different methods in Fig. 7 and Fig. 8. From Fig. 7, comparing with only using source domain method, our domain adaptation method exhibits improved recognition capability for overlapping and missed nuclei by incorporating an output of detection network in CPL. In Fig. 8, our results are closest to the nuclei segmentation ground truth. In contrast, SIFA[17] and PDAM[18] exhibit poorer segmentation performance on some overlapping nuclei. These results further demonstrate the effectiveness of our proposed method in domain adaptation, whether from pan-cancer to specific cancer types or from pan-cancer to other pan-cancer datasets.
Figure 8: The visualization comparisons between three domain adaptation methods and our proposed DAWN from PanNuke dataset to TNBC dataset.
TABLE VI: The ablation experiments of CFC and IST modules on TNBC, CryoNuSeg, Lizard and ConSeP datasets.
Data CFC IST DICE AJI DQ SQ PQ
0.8216 0.6764 0.8158 0.8094 0.6610
TNBC 0.8405 0.7045 0.8346 0.8128 0.6796
0.8340 0.7005 0.8236 0.8122 0.6699
0.8561 0.7280 0.8425 0.8202 0.6925
0.7857 0.4699 0.6121 0.7467 0.4667
CryoNuSeg 0.7946 0.5006 0.6279 0.7404 0.4668
0.7794 0.4734 0.5750 0.7479 0.4319
0.8041 0.5079 0.6374 0.7439 0.4760
0.6963 0.3579 0.4508 0.7316 0.3318
Lizard 0.7236 0.4081 0.5038 0.7344 0.3752
0.7022 0.3627 0.4582 0.7007 0.3476
0.7282 0.4031 0.5095 0.7364 0.3768
0.7813 0.4859 0.6101 0.7612 0.4644
ConSeP 0.7959 0.5010 0.6146 0.7621 0.4695
0.7899 0.4547 0.6138 0.7584 0.4675
0.8047 0.5093 0.6272 0.7591 0.4772
IV-F Ablation study
IV-F1 Evaluation of CFC and IST
To prove the effectiveness of CFC and IST components, we perform ablation studies with results listed in Table VI. On the TNBC and Lizard datasets, the five metrics have noticeable improvements whether CFC or IST is added. On CryoNuSeg and ConSeP datasets, when the CFC module is used, the DQ and PQ also obtain improvement, and the DQ and AJI decrease slightly. At the same time, when adding CFC and IST components, the model achieves the best performance, indicating that CFC and IST modules need to be used together. This phenomenon can be attributed to the significant morphological differences between the CryoNuSeg (ConSeP) and the source domain datasets. Only using CFC might lead to feature perturbations of the segmentation network. However, adding IST can guide the network in performing domain transfer toward the target domain. Our ablation study proves that the CFC and IST promote domain adaptation ability and enhance nuclei segmentation performance on weakly supervised scenarios.
TABLE VII: The mask fusion and distance filter ablation experiments on TNBC dataset.
Methods PanNuke $\rightarrow$ TNBC
DICE AJI DQ SQ PQ
w/o fusion 0.8418 0.7116 0.8318 0.8226 0.6845
w/o distance 0.8357 0.7045 0.8258 0.8232 0.6814
ours(d=10) 0.6541 0.4917 0.5771 0.7054 0.4170
ours(d=25) 0.8561 0.7280 0.8425 0.8202 0.6925
ours(d=40) 0.8332 0.7120 0.8313 0.8220 0.6850
Figure 9: The iteration ablation study on four datasets. The x-axis represents the number of iterations, and the y-axis represents model performance.
Figure 10: The iteration training loss on TNBC dataset. The x-axis represents the training step and the y-axis represent the training loss and F1 value.
TABLE VIII: The detection results before and after domain adaptation.
Datasets Stage Recall Precision F1
TNBC Bef. 0.7647 0.8949 0.8205
Aft. 0.8605 (0.10$\uparrow$) 0.7870 (0.11$\downarrow$) 0.8221 (0.002$\uparrow$)
CryoNuSeg Bef. 0.4637 0.7522 0.5737
Aft. 0.6635 (0.20$\uparrow$) 0.7326 (0.02$\downarrow$) 0.6963 (0.12$\uparrow$)
Lizard Bef. 0.7804 0.8373 0.7958
Aft. 0.8029 (0.02$\uparrow$) 0.8109 (0.03$\downarrow$) 0.8069 (0.01$\uparrow$)
ConSeP Bef. 0.5563 0.7893 0.6526
Aft. 0.7878 (0.23$\uparrow$) 0.5999 (0.19$\downarrow$) 0.6812 (0.03$\uparrow$)
Figure 11: The nuclei detection results. (a) represents the detection results of source only method; (b) represents the detection results of our DAWN.
IV-F2 Evaluation of Hyper-parameter in CPL
To prove that mask fusion and distance filter effectively improve segmentation performance, we perform an ablation study and parameter-sensitive experiments on the TNBC dataset. The results are shown in Table VII. From the result, we can see that, whether removing mask fusion or distance filter, the performance decreases evidently, which indicates the two components are effective in the CPL module because they can enhance the quality of pseudo-labels. Besides, we can see that the model is sensitive to the distance filter method. When the distance is small, the performance decreases because it destroys the nuclear shape. When the distance is enlarged, the performance is close to the result without a distance filter. Overall, the mask fusion and distance filter in CPL have also been proved to improve performance.
IV-F3 Evaluation of Iterative Training
We perform the experiments across the four datasets to evaluate the impact of the number of iterations in Fig. 9. In this figure, the x-axis represents the number of iterations, corresponding to the parameter ”r” in Algorithm 1, while the y-axis represents the values of the DICE, AJI, and PQ metrics. The results show that on the TNBC and Lizard datasets, the model’s performance shows an increasing trend as the number of training iterations increases. Specifically, the TNBC dataset achieves optimal performance at r=3, while the Lizard dataset performs best at r=4. In contrast, on the CryoNuSeg and ConSeP datasets, the model’s performance remains relatively stable after the second round of training. Consequently, we select the optimal performance according to r as the segmentation result. Furthermore, Fig. 10 presents the loss function and F1 curve for three iterations of training on the TNBC dataset, with each iteration consisting of 20 training epochs. From this figure, it can be observed that as training progresses, the model’s loss steadily decreases. Correspondingly, the F1 score increases with training and performs best in the final iteration.
IV-F4 Evaluation of Domain Adaptation
To evaluate the influence of our proposed auxiliary detection branch on nuclei detection ability, we conduct a domain adaptation comparison experiment on four target datasets in three metrics, i.e., Recall, Precision, and F1. The experimental results are shown in Table VIII. As seen from the table, when using the detection branch to assist domain transfer, the Recall has been significantly improved without reducing F1, which means that the auxiliary task can help the segmentation network to identify the nucleus more completely. This conclusion can also be verified from the visualization results in Fig. 11. The figure’s red $\star$ represents the actual nuclei location. Orange $\bullet$” and blue “+” represent the predicted location before and after domain adaptation. The figure shows that after adding the detection task, the predicted results are highly coincident with the actual labels, indicating that our detection task helps enhance the nuclei recognition ability of the segmentation task.
V Conclusion
To reduce the dependence of nuclei segmentation on manual annotation, we propose a domain-adaptive weakly supervised segmentation method named DAWN. Our method can transfer the model trained on the public data (source domain) to arbitrary external data (target domain) without fine pixel-level annotation. For this purpose, the domain adaptation of the segmentation network is guided by an auxiliary detection task. In DAWN, a consistent feature constraint (CFC) module is used to enhance the domain transfer efficiency at the feature level. In addition, a combined pseudo-labels optimization (CPL) method and interactive supervision training (IST) method are proposed to enhance the domain adaptation ability at the task level. The ablation experiments prove these components are effective in boosting the nuclei segmentation performance. Furthermore, our comparison experiments show that our DAWN outperforms existing weakly supervised methods and achieves similar performance compared with supervised methods.

### v3

@@
 Abstract
-Weakly supervised segmentation methods have gained significant attention due to their ability to reduce the reliance on costly pixel-level annotations during model training.
-However, the current weakly supervised nuclei segmentation approaches typically follow a two-stage pseudo-label generation and network training process.
-The performance of the nuclei segmentation heavily relies on the quality of the generated pseudo-labels, thereby limiting its effectiveness.
-This paper introduces a novel domain-adaptive weakly supervised nuclei segmentation framework using cross-task interaction strategies to overcome the challenge of pseudo-label generation.
-Specifically, we utilize weakly annotated data to train an auxiliary detection task, which assists the domain adaptation of the segmentation network.
-To enhance the efficiency of domain adaptation, we design a consistent feature constraint module integrating prior knowledge from the source domain.
-Furthermore, we develop pseudo-label optimization and interactive training methods to improve the domain transfer capability.
-To validate the effectiveness of our proposed method, we conduct extensive comparative and ablation experiments on six datasets.
-The results demonstrate the superiority of our approach over existing weakly supervised approaches.
-Remarkably, our method achieves comparable or even better performance than fully supervised methods.
-Our code will be released in https://github.com/zhangye-zoe/DAWN.
+Weakly supervised segmentation methods have garnered considerable attention due to their potential to alleviate the need for labor-intensive pixel-level annotations during model training.
+Traditional weakly supervised nuclei segmentation approaches typically involve a two-stage process: pseudo-label generation followed by network training.
+The performance of these methods is highly dependent on the quality of the generated pseudo-labels, which can limit their effectiveness.
+In this paper, we propose a novel domain-adaptive weakly supervised nuclei segmentation framework that addresses the challenge of pseudo-label generation through cross-task interaction strategies.
+Specifically, our approach leverages weakly annotated data to train an auxiliary detection task, which facilitates domain adaptation of the segmentation network.
+To improve the efficiency of domain adaptation, we introduce a consistent feature constraint module that integrates prior knowledge from the source domain.
+Additionally, we develop methods for pseudo-label optimization and interactive training to enhance domain transfer capabilities.
+We validate the effectiveness of our proposed method through extensive comparative and ablation experiments conducted on six datasets.
+The results demonstrate that our approach outperforms existing weakly supervised methods and achieves performance comparable to or exceeding that of fully supervised methods.
+Our code is available at https://github.com/zhangye-zoe/DAWN.
 Index Terms:
@@
 I Introduction
-Nuclei segmentation [1, 2]plays a vital role in computational pathology, serving as a fundamental step for various nuclei-related tasks.
-Accurate segmentation enables advanced analysis, such as nuclei classification [3, 4, 5], nuclei counting [6, 7] and nuclei tracing [8, 9, 10].
-However, developing supervised nuclei segmentation models in whole slide images (WSIs) is challenging due to the extensive requirement of precise pixel annotation in these images, which consist of gigapixels.
-At the same time, manual labeling of nucleus boundaries is time-consuming and demands domain expertise.
-To alleviate the manual burden, a weakly supervised segmentation algorithm is a promising solution, which leverages the inherent image structure and context to achieve segmentation, facilitating more efficient and accurate nuclei analysis.
+Nuclei segmentation plays a crucial role in computational pathology, serving as a fundamental step for various pathological diagnosis and patient prognosis tasks [1, 2, 3, 4].
+Meanwhile, accurate segmentation enables advanced nuclei analysis, including nuclei classification [5, 6, 7], nuclei counting [8, 9], and nuclei tracing [10, 11, 12].
+However, developing fully supervised nuclei segmentation models for whole slide images (WSIs) poses challenges because of the extensive requirement of precise pixel-wise annotations in these gigapixel-sized images.
+To ease the manual annotation burden, the weakly supervised segmentation algorithm provides a promising solution.
+It leverages weakly supervised annotation information to excavate the morphology and structure of objects, enabling more efficient and accurate nucleic analysis.
 Figure 1: The weakly supervised nuclei segmentation flows based on point annotation.
@@
 The point annotation is utilized to train a detection network, which can assist target domain transfer.
-Weakly supervised point annotation methods have recently been widely adopted for nuclei segmentation [11, 12, 13], which can decrease manual consumption by a large margin.
-However, due to inadequate supervision information, these methods generally perform a pre-processing step to generate masks for model training, as shown in Fig.
-1 (a).
-Subsequently, these generated pseudo-labels are utilized to guide the training of the segmentation network.
-However, this two-stage training approach still faces two significant challenges.
-Firstly, the pseudo-labels generated through pre-processing often exhibit a coarse granularity and lack precise representation of nuclear morphology.
+Weakly supervised methods based on point annotation have recently been widely adopted for nuclei segmentation [13, 14, 15], which reduce manual consumption by a large margin.
+Due to inadequate supervision information, these methods generally need a pre-processing step to generate masks.
+Then, these generated masks are utilized to guide the training of the segmentation network, as shown in 1 (a).
+However, the two-stage training flow faces two significant challenges.
+On the one hand, the generated pseudo-labels often are coarse and lack precise representation of nuclear morphology.
 For example, the Voronoi method [13] utilizes distance measurements to differentiate overlapping nuclei, which results in the disruption of boundary morphology, as depicted in Fig.
@@
 The level set method (LSM) [16] employs color features to simulate nuclear contours.
-However, it fails to distinguish tissues with similar colors, as illustrated in Fig.
+But, it fails to distinguish cytoplasm and nucleus with similar colors, as illustrated in Fig.
 2 (c).
-Secondly, these two-stage methods still have problems handling the suboptimal pseudo-labels.
-Although some single-stage optimization methods have been proposed to enhance segmentation performance, such as utilizing boundary mining techniques [15] to improve the quality of pseudo-labels and employing conditional random fields (CRF) [16] for post-processing to rectify segmentation results, these localized improvement methods fail to adapt to the entire training process, resulting in segmentation outcomes that remain unsatisfactory.
-Therefore, adopting a new training strategy is necessary to mitigate the impact of pseudo-label generation on nuclei segmentation results.
+On the other hand, these two-stage methods still suffer from the influences of inaccurate pseudo-labels.
+Although some localized optimization methods have been proposed to enhance segmentation performance, such as boundary mining [17] and conditional random fields (CRF) [18], these localized optimizing methods fail to adapt to the entire training process, resulting in segmentation outcomes that remain unsatisfactory.
+Therefore, it is necessary to adopt a new training paradigm to improve the nuclear segmentation performance in weakly supervised scenarios.
 Figure 2: The problems of generated pseudo labels.
 (a) represents the input image; (b) represents the generated mask using Voronoi [13]; (c) represents the generated mask using LSM [16]; (d) is the ground-truth label of input.
-Recent research has demonstrated the potential of domain adaptation in nuclei instance segmentation [17, 18, 19].
-They leverage prior information from the source domain and excavate domain-invariant features from the source data to achieve cross-domain task transfer, resulting in remarkable performance.
-Therefore, it is reasonable to believe that domain-adaptive methods can enhance instance segmentation performance in weakly supervised scenarios.
-On the one hand, domain adaptation methods do not explicitly generate pseudo-labels for supervising network training [20, 21].
-Instead, they enhance the network’s performance in the target domain through domain transfer strategies.
-It can mitigate the impact of inaccurate pseudo-labels during model training.
-On the other hand, compared to unsupervised domain adaptation methods, weakly supervised domain adaptation has evident superiority for providing additional supervision information.
-In the context of weak supervision, incorporating an auxiliary task, such as object detection, can leverage the weak annotations to effectively enhance the segmentation network’s ability to achieve domain transfer [22, 23, 24].
+Recent research has highlighted the potential of domain adaptation (DA) for improving segmentation tasks.
+By utilizing task-relevant information from the source domain, DA can improve instance segmentation performance in the target domain.
+However, existing unsupervised domain adaptation methods [19, 20, 21] often struggle with dense segmentation tasks like nuclei segmentation, primarily due to their limited ability to incorporate target domain information effectively.
+In weakly supervised scenarios, point-annotated images provide valuable prior knowledge regarding the location and distribution of nuclei.
+This additional information can be leveraged to enhance domain adaptation models.
+To address this, we propose a novel weakly supervised nuclei segmentation network incorporating domain adaptation techniques.
+Our approach presents several advantages over previous models.
+First, unlike traditional weakly supervised segmentation methods that rely heavily on generating pseudo-labels for training [22, 23], our method employs weak annotation to train auxiliary task to boost domain transfer capabilities.
+This approach minimizes the issues associated with inaccurate pseudo-labels.
+Second, our weakly supervised segmentation model provides distinct advantages over unsupervised methods by incorporating additional supervisory information.
+This integration allows us to leverage weak annotation information more effectively, significantly enhancing domain transfer capability [24, 25, 26].
 Motivated by these observations, this paper proposes a Domain-Adaptive Weakly supervised Nuclei segmentation method (DAWN).
-By utilizing point annotation data, DAWN incorporates a cross-task interaction strategy at both the feature and output levels, aiming to enhance the model’s domain transfer capability for accurate nuclei segmentation in weak supervision scenarios.
-Firstly, we propose a consistent feature constraint (CFC) module, which leverages the task correlation between the detection and segmentation networks to preserve prior knowledge of the target data.
+By utilizing point annotation data, DAWN incorporates a cross-task interaction strategy at both the feature and task levels to enhance the model’s domain transfer capability for accurate nuclei segmentation.
+Firstly, we propose a consistent feature constraint (CFC) module, which leverages the task correlation between the detection and segmentation networks to preserve prior knowledge related to the target data.
 Secondly, we introduce a task-level supervision information optimization strategy to address the challenge of the domain mismatch problem.
@@
   • We propose a novel domain-adaptive weakly supervised nuclei segmentation framework that departs from traditional two-stage weakly supervised methods.
-Our approach leverages a detection task with point annotations and domain transfer strategies to address the issue of inaccurate pseudo-label generation.
-  • We devise a consistent feature constraint module, which utilizes prior information of the source domain to improve the nuclei segmentation performance of target domain.
-  • We develop a combined pseudo-label optimization method incorporating the output information from the segmentation and the detection networks.
+Our approach leverages a detection task with point annotations and domain transfer strategies for accurate segmentation.
+  • We devise a consistent feature constraint module to retain the prior information from the source domain, which boosts the domain transfer efficiency at the feature level.
+  • We develop a combined pseudo-label optimization method incorporating the outputs from segmentation and detection networks.
 Besides, through an iterative training strategy, this module enhances the domain transfer capability at the task level.
-  • We propose an interactive supervised training method that leverages optimized pseudo-labels to simultaneously supervise the training of segmentation and detection networks.
+  • We propose an interactive supervised training method that leverages optimized pseudo-labels to supervise the training of segmentation and detection networks simultaneously.
 This approach facilitates the learning of complementary features between the two networks.
   • We validate our method on six datasets.
-Extensive experiments indicate that our method outperforms existing weakly supervised methods and achieves similar or even better performance with fully supervised methods.
+Extensive experiments indicate that our method outperforms existing weakly supervised methods and achieves equal or even better performance with fully supervised methods.
 II Related Work
-II-A Fully Supervised Nuclei Segmentation
-Nuclei segmentation is an essential workflow, and analysis based on nuclei can provide a basis for downstream tasks of computational pathology.
-In the past, watersheds [25] or Otsu threshold[26] usually led to over-segmentation or under-segmentation problems.
-With the successes of U-Net[8] in the cell tracing task, the variations of U-Net are emerging [27, 28, 29] and widely used in the nuclei segmentation.
-At present, nuclei segmentation methods mainly focus on solving the nuclei overlapping and unclear staining problems.
-To solve nuclei overlapping, HoverNet[30] calculates the horizontal and vertical gradients to supervise training.
-Cellpose[31] uses simulated diffusion and spatial gradients to trace instances to which each pixel belongs.
-CDNet[32] employs the direction difference map to distinguish different instances of overlapping regions.
-Besides, DCAN [33] adds the contour category to the semantic prediction task to separate instances.
-To solve the unclear staining problem, some staining separation methods [34, 35] are proposed to extract nuclei staining components.
-In addition, Sams-net [36] and Triple-Unet [37] use H-channel of staining separation images to enhance the nuclei feature representation ability.
-Although these methods achieve good progress in nuclei segmentation, they need a large amount of pixel-level manual annotation, which is expensive and labor-intensive.
-On the contrary, weakly supervised methods provide an effective solution to reduce the reliance of model development on annotations.
-They allow training models with coarse annotations, reducing time and manual costs.
-II-B Weakly Supervised Segmentation
-Existing weakly supervised segmentation labels mainly include image-level, bounding-box, and point annotations.
-Among that, the models using image-level annotations typically rely on the Class Activation Map (CAM) [38] to generate pseudo labels.
-PRM[39] utilizes peak backpropagation from CAMs to generate highly informative object regions.
-SC-CAM[40] accelerates the integrity of CAMs using the unsupervised clustering sub-category information.
-RRM[41] uses the CAM and reliable score evaluation to generate finer pseudo labels.
-The methods utilizing bounding-box annotations mainly focus on the saliency models.
-BoxCaseg[42] employs the salient segmentation head to enhance the object localization ability.
-BoxTeacher[43] uses noise-aware pixel loss and noise-reduced affinity loss to optimize the pseudo labels.
-Due to the small nuclei morphology, point labels are widely adopted as the supervision information in pathological images.
-WeakSeg [16] develops Voronoi splitting and clustering to generate pseudo labels.
-After that, Voronoi is widely used to generate segmentation masks[11, 44, 45, 13].
-Besides, [46] establishes a propagating detection map to segment fluorescence images.
-[12] adopts a co-training strategy to segment phase images.
-Even though these weakly supervised nuclei segmentation methods have achieved good results, they often lead to poor shape prediction due to the need for more morphological consideration.
-In this paper, we improve the nuclear morphology through a pre-training model and domain adaptation strategies.
-II-C Domain Adaptation in Medical Image Analysis
-Pre-training methods [47] can bring convenience to small-sample task training.
-However, directly using pre-trained models often leads to poor performance due to the distribution differences of samples.
-Therefore, some domain adaptation methods are proposed to address the problem of domain transfer.
-[48] uses the image translation method to align the pixel value distribution, reducing the gap between the source domain and the target domain.
-DIGA[49] develops a distribution adaptation and a semantic adaptation module to enable model adaptation jointly.
-Inspired by natural scenario methods[50], domain adaptation is also used in medical image analysis[17].
-DARCNN[19] implements domain adaptation from natural scene images to hematoxylin-eosin (H&E) images, exploiting domain-invariant and domain-specific features.
-CyC-PDAM[51] transfers the model from the fluorescence images domain to the H&E images domain.
-CAPL-Net [52] implements domain adaptation in H&E images from different sources.
-SAMDA [53] combines the Segment Anything Model (SAM) with nnUNet [54] in the embedding space to achieve a few-shot domain adaptation.
-However, existing models mainly utilize unsupervised methods for domain adaptation, often leading to poor performance due to the lack of understanding of target domain features.
-In this paper, we leverage auxiliary tasks to establish connections between the source domain model and the detection task, which promotes the transfer efficiency of the source model to the target domain.
+II-A Nuclei Instance Segmentation
+Pathological image analysis based on nuclear segmentation [27, 28, 29, 30, 31] has been widely discussed.
+Currently, some representative work mainly focuses on the problem of nuclear boundary overlapping.
+Such as DCAN [32] adds the contour prediction task to the semantic segmentation model to separate instances.
+HoverNet[33] uses horizontal and vertical gradients to distinguish overlapping instances.
+Cellpose[34] uses simulated diffusion and spatial gradients to trace the instances to which each pixel belongs.
+In addition, other methods [35, 36] are also proposed to solve the boundary overlapping problem.
+Although fully supervised nuclei segmentation has achieved good progress, it is expensive and labor intensive because it requires a lot of manual annotation at the pixel level.
+In contrast, weakly supervised methods provide an effective solution to reduce the reliance of model development on annotations.
+Mainstream weakly supervised nuclear segmentation models follow a two-stage training paradigm.
+First, point annotation images are used to generate pseudo-labels, and then these generated pseudo-labels are used to supervise network training [13, 37, 38, 39].
+However, these two-stage training methods are affected by initialized pseudo-labels, and there is no explicit optimization in the training process, so the model performance is poor.
+In this paper, we propose a domain adaptation method for weakly supervised nuclear segmentation.
+Our method uses end-to-end training to avoid the traditional two-stage training paradigm and significantly improves the accuracy of nuclear segmentation in weakly supervised scenarios.
+II-B Weakly Supervised Segmentation Methods
+Existing weakly supervised segmentation approaches utilize various types of partially labeled annotation to help automatically pixel-level segmentation.
+These weak labels include image-level, box-level, and point labels.
+Among these, methods relying on image-level annotations typically use class activation maps (CAMs) [40] to generate pseudo-labels.
+For instance, PRM [41] employs CAM peak backpropagation to identify highly informative object regions.
+SC-CAM [42] enhances CAM completeness by incorporating information from unsupervised clustering.
+RRM [43] combines CAM with a reliable score evaluation to produce more refined pseudo-labels.
+In addition, several other models [44, 45] are proposed to solve weakly supervised video segmentation tasks.
+The box-level annotations are often used in conjunction with saliency models.
+BoxCaseg [46] leverages a salient segmentation head to improve object localization, while BoxTeacher [47] applies noise-aware pixel loss and noise-reduced affinity loss to optimize pseudo-labels.
+Point labels [48, 31, 49] are widely used in small object segmentation tasks because of their excellent location ability.
+ASDT [39] propagates labels between image and pixel levels, thereby improving semantic segmentation performance.
+Despite the success of existing weakly supervised segmentation methods, these approaches often struggle with accurate shape prediction due to insufficient consideration of the morphological characteristics of instances.
+In this paper, we introduce the domain adaptive method into the weakly supervised segmentation model and learn morphological features with the prior knowledge of external data, thus improving the segmentation performance.
+II-C Domain Adaptation Segmentation
+Pre-training methods [50] have made it possible to train models with limited samples.
+However, directly applying pre-trained models to target domains often fails due to distribution gaps between source and target samples.
+This challenge has led to the development of domain adaptation techniques.
+For example, Wang et al.
+[51] use image translation methods to align pixel value distributions, thereby reducing the domain gap.
+Similarly, DIGA [52] introduces both distribution adaptation and semantic adaptation modules to facilitate joint model adaptation.
+ADPL [53] improves domain adaptation by constructing an adaptive dual-path network that bridges source and target domains.
+CyCADA [54] combines spatial alignment of generated images with potential representation alignment to achieve unsupervised domain adaptation.
+In a different approach, DDMRL [55] addresses object detection by employing two stages domain diversity and multi-domain invariant representation learning.
+WDASS [56] proposes a weakly-supervised domain adaptive semantic segmentation method that leverages prototypical contrastive learning to align features across domains.
+Drawing inspiration from natural scene methods [57, 58, 59], domain adaptation has also been applied to medical image analysis [19].
+DARCNN [21] adapts models from natural scene images to hematoxylin-eosin (H&E) stained images, capturing both domain-invariant and domain-specific features.
+CyC-PDAM [60] transfers models between fluorescence images and H&E images, while CAPL-Net [61] focuses on domain adaptation for H&E images from different sources.
+SAMDA [62] integrates the Segment Anything Model (SAM) with nnUNet [63] in the embedding space to enable few-shot domain adaptation.
+Despite the progress made by these domain adaptation methods, their performance often suffers due to insufficient understanding of target domain characteristics.
+In this paper, we propose leveraging auxiliary tasks to bridge the gap between the source domain model and the detection task, enhancing the efficiency of model transfer to the target domain.
 Figure 3: The schematics of our proposed DAWN framework.
@@
 III-A Framework Overview
-Detection and segmentation are closely related tasks that share similar semantic information, such as the location and texture of the object, which can be utilized to assist in instance detection or segmentation.
-In multiple-task learning scenarios, detection and segmentation networks are usually combined to improve the performance of related tasks [55, 56].
-At that point, under the domain adaptation background, we design a detection task based on point annotations to assist the segmentation network domain transfer.
-By enhancing the adaptability of the segmentation network to target domain data, the nuclei segmentation accuracy can be improved in weakly supervised scenarios.
-This paper establishes cross-task interaction strategies at both the feature and task levels.
-The feature interaction is designed to preserve the prior segmentation information related to the target domain.
-At the same time, the task interaction is used to compensate for the shortcomings of the segmentation output through combined pseudo-label optimization and interactive supervision methods.
-Overall, the flowchart of our proposed DAWN is depicted in Fig.
+Detection and segmentation are closely related tasks that share semantic information, such as object location and texture, which can be leveraged to enhance both object detection and instance segmentation.
+In multi-task learning, detection and segmentation networks are often combined to improve the performance of both tasks [64].
+In the context of domain adaptation, we design a detection task using point annotations to support the segmentation network’s adaptation to the target domain.
+By improving the segmentation network’s adaptability to target domain data, this approach enhances nuclei segmentation accuracy, particularly in weakly supervised scenarios.
+This paper introduces cross-task interaction strategies at both the feature and task levels.
+The feature-level interaction is designed to retain prior segmentation information relevant to the target domain.
+In contrast, task-level interaction compensates for the limitations of segmentation output through a combination of pseudo-label optimization and interactive supervision techniques.
+The overall flow of our proposed DAWN is illustrated in Fig.
 3.
-We firstly pre-train a segmentation network in source domain as shown in Fig.
+We first pre-train the segmentation network in the source domain, as illustrated in Fig.
 3 (a).
-To overcome the domain shift problem caused by data distribution variations, we design a consistent feature constraint (CFC) module and a combined pseudo-label optimization (CPL) module as shown in Fig.
+To address the domain shift caused by variations in the data distribution, we introduce a Consistent Feature Constraint (CFC) module and a Combined Pseudo-Label (CPL) optimization module, as shown in Fig.
 3 (c)-(d).
-In addition, to promote domain transfer ability and efficiency, we employ an interactive supervision training (IST) method.
-According to these constraints, our training loss function is formulated as follows:
-$Loss=L_{det}+\alpha L_{fea}+\beta L_{dyn},$ (1)
-where $L_{det}$ is the loss of detection network, $L_{fea}$ is the consistent feature constraint loss, and $L_{dyn}$ is the dynamic supervision loss.
-$\alpha$ and $\beta$ are used to balance the weight of each loss.
-It should be noted that $L_{det}$ corresponds to a fully-supervised learning task.
+Additionally, to enhance the efficiency and effectiveness of domain transfer, we employ an Interactive Supervision Training (IST) method.
+Based on these components, our overall training loss function is defined as follows:
+$Loss=L_{\text{det}}+\alpha L_{\text{fea}}+\beta L_{\text{dyn}},$ (1)
+where $L_{\text{det}}$ represents the detection network loss, $L_{\text{fea}}$ is the consistent feature constraint loss, and $L_{\text{dyn}}$ corresponds to the dynamic supervision loss.
+The parameters $\alpha$ and $\beta$ are used to balance the weight of different loss terms.
+It should be noted that $L_{\text{det}}$ is associated with a fully supervised learning task.
 III-B Consistent Feature Constraint
@@
 3 (a).
-In this step, the loss function is formulated as follows:
-$\begin{split}L_{pre}&=CE(\hat{p}_{s},p_{s})+Dice(\hat{p}_{s},p_{s})\\ &+MSE(\hat{h}_{s,x},h_{s,x})+MSE(\hat{h}_{s,y},h_{s,y}),\end{split}$ (2)
-where $\hat{p}_{s}$, $\hat{h}_{s,x}$, and $\hat{h}_{s,y}$ represent the foreground probability prediction, horizontal distance prediction, and vertical distance prediction.
+In this step, the loss function [33] is formulated as follows:
+$\begin{split}L_{pre}&=\text{CE}(\hat{p}_{s},p_{s})+\text{Dice}(\hat{p}_{s},p_{s})\\ &+\text{MSE}(\hat{h}_{s,x},h_{s,x})+\text{MSE}(\hat{h}_{s,y},h_{s,y})\\ &+\text{MSE}(\nabla\hat{h}_{s,x},\nabla h_{s,x})+\text{MSE}(\nabla\hat{h}_{s,y},\nabla h_{s,y}),\end{split}$ (2)
+where $\hat{p}_{s}$, $\hat{h}_{s,x}$, and $\hat{h}_{s,y}$ represent the predicted foreground probability, horizontal distance, and vertical distance.
 CE, Dice, and MSE represent cross-entropy loss, dice loss, and mean square error loss.
-In the following domain adaptation process, we jointly train the segmentation and detection networks, and the process is shown in Fig.
+During the domain adaptation stage, we jointly train the segmentation and detection networks, and the process is shown in Fig.
 3 (c) and Stage 2 of Algorithm 1.
-Firstly, the well-trained segmentation network is utilized to initialize the target domain segmentation model.
-Then, we map the output feature $f_{seg}$ of encoder into $f_{seg}^{c}$ by a convolution layer and a fully connection layer:
+First, the well-trained segmentation network is utilized to initialize the target domain segmentation model.
+And the encoder feature of the segmentation network is denoted as $f_{seg}$.
+Then, we map the feature $f_{seg}$ into a new feature space by a convolution layer and a fully connected layer and the output feature is denoted as $f_{seg}^{c}$ as follows:
 $f_{seg}^{c}=FC(Conv(f_{seg})),$ (3)
-where $FC$ and $Conv$ represent fully connected and convolution layers, respectively.
-Finally, the output feature $f_{seg}^{c}$ is used to compute the consistent feature constraint.
-It should be noted that the detection network is trained in the target domain, and the segmentation network is initialized in the source domain.
-Hence, the CFC can force the segmentation network to gradually close to the target domain.
-Considering single central pixel of nuclei just represents nuclei position while ignoring nuclei texture, it is hard to train an effective detection network using point label $y_{p}^{t}$ [57, 58].
-Thus, we employ extended Gaussian distribution [16] to model nuclei morphology.
+where $FC$ and $Conv$ represent fully connected and convolution layers. Finally, the output feature $f_{seg}^{c}$ is used to compute the consistent feature constraint.
+In the detection network, considering single central pixel just represents nuclei position while ignoring nuclei texture, it is hard to train an effective detection network using target label $y_{t}^{p}$ [65].
+Thus, we employ an extended Gaussian distribution [18] to model the morphology of nuclei.
 The extended Gaussian encoding process is shown in Fig.
-3 (b) and can be formulated as follows:
+3 (b) and formulated as follows:
 $\begin{split}M_{g,i}=\left\{\begin{array}[]{ll}exp(-\frac{{D_{i}}^{2}}{2{\sigma}^{2}}),&\;D_{i}\leq r_{1},\\ 0,&\;r_{1}<D_{i}\leq r_{2},\\ -1,&otherwise,\end{array}\right.\end{split}$ (4)
-where $D_{i}$ represents the distance between pixel $i$ and nearest point annotation, $r_{i}$ represents initialized nuclei radius, the region between $r_{1}$ and $r_{2}$ is background, $\sigma$ is the discrete degree of Gaussian distribution.
+where $D_{i}$ represents the distance between pixel $i$ and nearest point annotation, $r_{1}$ represents initialized nuclei radius, the region between $r_{1}$ and $r_{2}$ is background, $\sigma$ is the discrete degree of Gaussian distribution.
 In $M_{g,i}$, the pixel corresponding to $-1$ does not participate in the training.
 Employing the generated encoding, a supervised detection model is trained as follows:
-$L_{det}=\frac{1}{|\Omega|}\sum_{i\in\Omega}w_{i}(\hat{q}_{i}-M_{g,i})^{2},$ (5)
+$L_{det}=\frac{1}{|\Omega|}\sum_{i\in\Omega}w_{i}(\hat{q}_{t,i}^{(0)}-M_{g,i})^{2},$ (5)
 where $\Omega$ represents the pixel set participating training, i.e.
 $\Omega=\{i|0<D_{i}\leq r_{2}\}$; $w_{i}$ is the training weight of pixel $i$.
-In order to pay more attention to the foreground, we set $w_{i}$ equals $10$ when $D_{i}<r_{1}$; otherwise, $w_{i}$ equals $1$.
-The extended Gaussian encoding not only encodes the tissue texture but can also learn nuclei shape gradually, assisting segmentation network optimization.
-Similar to the above segmentation network, the middle layer output $f_{det}$ of the detection network is mapped into $f_{det}^{c}$ to calculate CFC as shown in Fig.
+In order to pay more attention to the foreground, we set $w_{i}$ equal to $10$ when $D_{i}<r_{1}$; otherwise, $w_{i}$ equals $1$.
+The extended Gaussian encoding not only encodes tissue texture, but can also learn nuclei shape gradually, assisting in segmentation network optimization.
+Similar to the above segmentation network, the encoder output $f_{det}$ of the detection network is mapped into new representation $f_{det}^{c}$, which will be used to calculate CFC with $f_{seg}^{c}$ as shown in Fig.
 3 (c) and the feature constraint loss $L_{fea}$ is calculated as follows:
@@
 where $dim$ represents the dimension of feature encoding.
-$f_{seg}^{c}$ and $f_{det}^{c}$ are mapped features of segmentation and detection networks.
-In this process, the CFC module forces the two networks to excavate the essential task correlation features, enhancing domain adaptation efficiency.
+$f_{seg}^{c}$ and $f_{det}^{c}$ are feature embeddings of segmentation and detection networks.
+Due to the segmentation model being trained on the source domain dataset, its feature representation is closely related to the source domain.
+In contrast, the detection model is trained on the target domain, so its features are more relevant to the target domain data.
+Leveraging the target domain representation of $f_{det}^{c}$, the CFC effectively aligns $f_{seg}^{c}$ with $f_{det}^{c}$, enhancing the domain adaptation process.
 Algorithm 1 The Whole Training Process of Our DAWN
@@
 10:  Input: Target domain data $D_{t}$; Point annotation $y_{t}^{p}$; initialized weight $W_{s}$;
-11:  Initialize: $r=4$; $e_{t}=20$; Initialize segmentation network with $W_{s}$;
+11:  Initialize: $I=4$; $e_{t}=20$; Initialize segmentation network with $W_{s}$;
 12:  Output: Instance segmentation results $\hat{y}_{t}$;
-13:  for iter=1, …, $r$ do
+13:  for iter=1, …, $I$ do
 14:     for epoch=1,…,$e_{t}$ do
-15:        Calculate loss: $L_{det}+\alpha L_{fea}+\beta L_{dyn}$ in Eq. 1 and update network;
+15:        Calculate loss: $L_{\text{det}}+\alpha L_{\text{fea}}+\beta L_{\text{dyn}}$ in Eq. 1 and update network;
 16:     end for
-17:     Prediction outputs: $\hat{p}_{t}^{seg}$, $\hat{q}_{t}^{det}$;
-18:     Update pseudo-label: $M_{pse}$leftarrow CPL($\hat{p}_{t}^{seg}$, $\hat{q}_{t}^{det}$, $y_{t}^{p}$);
+17:     Prediction outputs: $\hat{p}_{t}$, $\hat{q}_{t}$;
+18:     Update pseudo-label: $M_{pse}$leftarrow CPL($\hat{p}_{t}$, $\hat{q}_{t}$, $y_{t}$);
 19:     Training networks with $M_{pse}$ in Eq. 9;
@@
 III-C Combined Pseudo-label Optimization
-Due to domain shift resulting from nuclei morphology and size variations across different organs, there may be an under-segmentation problem in the target domain prediction $\hat{y}_{t}^{seg}$.
-To solve the problem, we devise a combined pseudo-label (CPL) optimization method by combining the outputs of the segmentation and detection networks.
-This method utilizes the superior object localization capabilities of the detection network and effectively compensates for the flaw of the segmentation network.
-The schematic diagram of the method is depicted in Fig.
+Due to domain shifts caused by variations in nuclei morphology and size across different organs, under-segmentation can occur in the target domain predictions.
+To address this issue, we propose a Combined Pseudo-Label (CPL) optimization method, which integrates the outputs of both the segmentation and detection networks.
+This approach utilizes the superior object location capabilities of the detection network to effectively compensate for the shortcomings of the segmentation network.
+A schematic diagram of the process is provided in Fig.
 4.
-The CPL method comprises two key components: mask fusion and distance filter.
+The CPL method comprises two key components: mask fusion and the distance filter.
 In mask fusion, the prediction outputs from both networks are combined to achieve comprehensive nuclei segmentation.
 The distance filter is then applied to refine the fused results by eliminating incorrectly segmented nuclei.
-The resulting outputs from the CPL serve as the ground truth annotations for supervising the subsequent training.
+The resulting outputs from the CPL serve as ground-truth annotations for supervising the subsequent training.
 III-C1 Mask Fusion
-Based on the prediction output $\hat{q}_{t}^{det}$ of the detection network, we employ a threshold filter to preserve the high confidence region of prediction $\hat{q}_{t}^{det}$ as shown in Fig.
+Based on the prediction output $\hat{q}_{t}$ of the detection network, we employ a threshold filter to preserve the high confidence region of prediction $\hat{q}_{t}$ as shown in Fig.
 4 (a) and obtain the initialized detection $M_{det}$ as shown in Fig.
 4 (b), which represents the nuclei region:
-$M_{det}=I_{[\hat{q}_{t}^{det}>\theta]}\in\{0,1\}^{N\times N},$ (7)
-where $\hat{q}_{t}^{det}$ represents the probability output of detection model, $\theta$ is the threshold value, $M_{det}$ represents generated nuclei region.
-At the same time, we transform the probability output $\hat{p}_{t}^{seg}$ of the segmentation network, obtaining the initialized $M_{ini}$, which has the under-segmentation problem as shown in Fig.
+$M_{det}=I_{[\hat{q}_{t}>\theta]}\in\{0,1\}^{N\times N},$ (7)
+where $\hat{q}_{t}$ represents the probability output of detection model, $\theta$ is the threshold value, $M_{det}$ represents generated nuclei region.
+At the same time, we transform the probability output $\hat{p}_{t}$ of the segmentation network, obtaining the initialized $M_{ini}$, which has the under-segmentation problem, as shown in Fig.
 4 (c).
@@
 III-D Interactive Supervision Training Method
-This sub-section proposes a dynamic pseudo-labels optimization method to facilitate domain adaptation in the segmentation network.
-Specifically, the pseudo-labels generated from the CPL module in the $(i-1)^{th}$ round will be viewed as supervision information for the training of segmentation and detection networks in the $i^{th}$ round.
-By continuously optimizing the segmentation network, this method helps the progressive domain transfer of the network.
-Our interactive supervision process is shown in Fig.
+This subsection proposes a dynamic pseudo-label optimization method to facilitate domain adaptation in the segmentation network.
+Specifically, the pseudo-labels generated from the CPL module in the $(t-1)^{th}$ round serve as supervision information for the training of segmentation and detection networks in the $t^{th}$ round.
+By iterative optimizing the segmentation network, this method enables progressive domain transfer.
+The interactive supervision process is shown in Fig.
 3 (c)-(d) and Algorithm 1 Line 13-16.
@@
 Thus, this method can be considered as an interactive supervision process, and the loss function for this method is shown below:
-$L_{dyn}=CE(M_{pse},\hat{p}_{t}^{seg})+MSE(M_{pse},\hat{q}_{t}^{det}),$ (9)
-where $M_{pse}$ is the pseudo-label generated by the CPL module, $\hat{p}_{t}^{seg}$ and $\hat{q}_{t}^{det}$ represent the predicted output of segmentation and detection networks in the target domain.
+$L_{dyn}=\text{CE}(M_{pse},\hat{p}_{t})+\text{MSE}(M_{pse},\hat{q}_{t}),$ (9)
+where $M_{pse}$ is the pseudo-label generated by the CPL module, $\hat{p}_{t}$ and $\hat{q}_{t}$ represent the predicted output of segmentation and detection networks in the target domain.
 By using the pseudo-labels to supervise the network training in a dynamic and progressive manner, our approach promotes the effectiveness of domain adaptation and enhances the performance of the segmentation network.
@@
 IV-A Datasets
-In this subsection, we briefly introduce the PanNuke, MoNuSeg used in the source domain (♣), and TNBC, CryoNuSeg, ConSeP used in the target domain (♠) in this paper.
-The PanNuke, MoNuSeg, TNBC, CryoNuSeg, and ConSeP datasets are sampled at ×40 magnification.
+In this subsection, we introduce the PanNuke, MoNuSeg datasets used in the source domain (♣), and TNBC, CryoNuSeg, ConSeP datasets used in the target domain (♠).
+In detail, the PanNuke, MoNuSeg, TNBC, CryoNuSeg, and ConSeP datasets are sampled at ×40 magnification.
 The Lizard dataset is sampled at ×20 magnification.
@@
 Meantime, a ResNet encoder combined with a DenseNet decoder is used as a segmentation network.
-Besides, all experiments are run with PyTorch on an Nvidia RTX 3090 GPU.
-We set loss weights $\alpha=0.1$, $\beta=0.15$ used in Eq.
-1.
-The hyperparameters used Eq.
+All experiments are run with PyTorch on an Nvidia RTX 3090 GPU.
+We set the loss weights in Eq.
+1 to $\alpha=0.1$ and $\beta=0.15$.
+The hyperparameters specified in Eq.
 4 and Eq.
-7 are shown in Table I.
-$r_{1}$ is set according to the mean nuclei radius of the validation set, and $r_{2}$ is set to double $r_{1}$ for balancing foreground and background pixels.
-$\theta$ is set according to the detection network output on the validation set.
-We evaluate our experiments over metrics of DICE, AJI[60], DQ[65], SQ[65] and PQ[65].
-Among those, AJI, DQ, SQ, and PQ are instance-level metrics.
-DQ evaluates the detection quality, SQ evaluates segmentation quality, and PQ is the product of the DQ and SQ.
-Throughout all the tables in this paper, we bold the best and underline the second best.
+7 are detailed in Table I.
+Specifically, $r_{1}$ and $\sigma$ are determined based on the mean and standard deviation of the nuclear radius in the validation set, with $r_{2}$ set at twice the value of $r_{1}$ to balance the foreground and background pixels.
+The parameter $\theta$ is established according to the output of the detection network in the validation set, while $d$ is determined empirically.
+We evaluate our experiments using the metrics of DICE, AJI[67], DQ[72], SQ[72] and PQ[72].
 TABLE II: The comparison of different weakly supervised and fully supervised methods.
   Datasets Methods DICE AJI DQ SQ PQ Datasets Methods DICE AJI DQ SQ PQ
+  C2FNet[15] 0.754 0.605 0.763 0.682 0.505 C2FNet[15] 0.597 0.213 0.286 0.600 0.163
 DDTNet[16] 0.789 0.665 0.807 0.745 0.601 DDTNet[16] 0.602 0.236 0.350 0.619 0.259
@@
 MaskGANet[38] 0.751 0.597 0.744 0.612 0.454 MaskGANet[38] 0.603 0.197 0.256 0.584 0.134
-C2FNet[13] 0.754 0.605 0.763 0.682 0.505 C2FNet[13] 0.597 0.213 0.286 0.600 0.163
-TNBC PseEdgeNet[66] 0.682 0.497 0.688 0.547 0.394 Lizard PseEdgeNet [66] 0.545 0.140 0.229 0.534 0.105
-(a) WeakSAM[67] 0.834 0.697 0.795 0.782 0.620 (c) WeakSAM[67] 0.643 0.279 0.314 0.608 0.267
+TNBC SPF-TN [44] 0.806 0.687 0.812 0.739 0.611 Lizard SPF-TN[44] 0.610 0.221 0.314 0.602 0.261
+(a) PENet [73] 0.682 0.497 0.688 0.547 0.394 (c) PENet [73] 0.545 0.140 0.229 0.534 0.105
+WeakSAM[74] 0.834 0.697 0.795 0.782 0.620 WeakSAM[74] 0.643 0.279 0.314 0.608 0.267
 U-Net [10] 0.783 0.692 0.754 0.750 0.617 U-Net [10] 0.716 0.365 0.436 0.633 0.353
@@
 Micro-Net [75] 0.813 0.673 0.772 0.795 0.615 Micro-Net [75] 0.776 0.519 0.615 0.759 0.451
+AOST [76] 0.821 0.694 0.822 0.814 0.674 AOST[76] 0.709 0.394 0.523 0.712 0.364
 DAWN (Ours) 0.856 0.728 0.843 0.820 0.693 DAWN (Ours) 0.728 0.403 0.510 0.736 0.377
   Datasets Methods DICE AJI DQ SQ PQ Datasets Methods DICE AJI DQ SQ PQ
+  C2FNet[15] 0.643 0.384 0.581 0.573 0.332 C2FNet[15] 0.749 0.464 0.497 0.706 0.398
 DDTNet[16] 0.696 0.440 0.618 0.668 0.389 DDTNet[16] 0.598 0.295 0.282 0.452 0.186
@@
 MaskGANet[38] 0.647 0.393 0.595 0.654 0.380 MaskGANet[38] 0.653 0.367 0.401 0.696 0.374
-C2FNet[13] 0.643 0.384 0.581 0.573 0.332 C2FNet[13] 0.749 0.464 0.497 0.706 0.398
-CryoNuSeg PseEdgeNet[66] 0.620 0.321 0.507 0.529 0.306 ConSeP PseEdgeNet [66] 0.569 0.226 0.245 0.307 0.171
-(b) WeakSAM[67] 0.754 0.480 0.627 0.685 0.449 (d) WeakSAM[67] 0.669 0.378 0.384 0.680 0.401
+CryoNuSeg SPF-TN [44] 0.676 0.449 0.611 0.634 0.391 ConSeP SPF-TN[44] 0.632 0.392 0.373 0.667 0.382
+(b) PENet[73] 0.620 0.321 0.507 0.529 0.306 (d) PENet [73] 0.569 0.226 0.245 0.307 0.171
+WeakSAM[74] 0.754 0.480 0.627 0.685 0.449 WeakSAM[74] 0.669 0.378 0.384 0.680 0.401
 U-Net [10] 0.697 0.469 0.546 0.731 0.403 U-Net [10] 0.761 0.499 0.569 0.725 0.434
@@
 Dist [68] 0.792 0.518 0.595 0.721 0.432 Dist[68] 0.793 0.512 0.587 0.716 0.422
+AOST [76] 0.754 0.487 0.602 0.724 0.459 AOST[76] 0.764 0.478 0.565 0.704 0.420
 DAWN (Ours) 0.804 0.508 0.637 0.744 0.476 DAWN (Ours) 0.805 0.509 0.627 0.759 0.477
 IV-C Comparison with Weakly Supervised Methods
-We compare our proposed domain-adaptive weakly supervised nuclei segmentation method with seven other models on four datasets, and the comparison methods consist of DDTNet[14], BoNuS[15], WeakSeg[16], MaskGANet[45], C2FNet[13], PseEdgeNet[66], and WeakSAM[67].
+We compare our proposed domain-adaptive weakly supervised nuclei segmentation method with eight other models on four datasets.
+The comparison methods consist of DDTNet[16], BoNuS[17], WeakSeg[18], MaskGANet[38], C2FNet[15], PENet [73], SPF-TN [44], and WeakSAM[74].
 The experiment results are shown in Table II.
-In the table, the best weakly supervised method is bold with blue text, and the second best weakly supervised method is underlined with blue text.
-For fair comparison, for the two-stage methods WeakSeg, BoNuS and DDTNet, we use the same pseudo labels as our DAWN.
-In detail, the pre-trained segmentation network is test on target dataset and the prediction outputs is used as pseudo labels.
-In the experiment, we use PanNuke as source domain dataset.
+In the table, the best weakly supervised method is bold, and the second best weakly supervised method is underlined.
+For a fair comparison, we use the same pseudo labels as our DAWN for the two-stage methods WeakSeg, BoNuS, and DDTNet.
+In detail, the pre-trained segmentation network is tested on the target dataset, and the prediction outputs are used as pseudo labels.
+In the experiment, we use PanNuke as the source domain dataset.
 IV-C1 Comparisons on Breast Cancer TNBC
-We first validate our proposed domain-adaptive weakly supervised nuclei segmentation on a breast cancer dataset, and the experimental results are shown in Table II.
+We validate our proposed DAWN on a breast cancer dataset, and the experimental results are shown in Table II (a).
 From this table, we observe that our proposed method outperforms existing state-of-the-art weakly supervised segmentation methods.
@@
 Moreover, our method exhibits significant advantages over traditional two-stage weakly supervised models such as MaskGANet and WeakSeg.
-Our method surpasses MaskGANet by at least 10% and WeakSeg by at least 4.0% on all five metrics.
+In detail, our method surpasses MaskGANet by at least 10% and surpasses WeakSeg by at least 4.0% on all five metrics.
 Furthermore, we conduct visualization comparison experiments as shown in the first two rows of Fig.
 5.
-From the figure, it can be observed that our proposed DAWN, WeakSeg, DDTNet, and BoNuS can effectively segment nuclei.
-However, methods like MaskGANet and PENet exhibit weaker morphological learning capabilities.
-We attribute this to pre-trained segmentation networks, where the original domain’s segmentation network contains a wealth of prior morphological information that improves nuclei segmentation performance.
+From the segmentation results, it can be seen that our method has clear advantages.
+On the TNBC and CryoNuSeg datasets, our method segments cell nuclei contours more accurately, while on the Lizard and ConSeP datasets, when the nuclei distribution is dense, our method can more comprehensively identify nuclei.
 IV-C2 Comparisons on Pan-Cancer CryoNuSeg
-We evaluate the performance of our proposed domain adaptation method for weakly supervised segmentation on a challenging pan-cancer dataset called CryoNuSeg.
-It is worth noting that the CryoNuSeg dataset exhibits significant variations in nuclei morphology, making it a problematic dataset for training.
-Our experimental results are presented in Table II (b).
-From these results, it can be observed that our proposed DAWN outperforms the best-performing method, WeakSAM, by 5.0% and 2.8% in terms of DICE and AJI metrics, respectively.
-Furthermore, compared to the second-best method, DDTNet, our DAWN achieves approximately 10% improvement in DICE, AJI, and PQ metrics.
+We evaluate the performance of our proposed method, DAWN, on the challenging pan-cancer CryoNuSeg dataset.
+This dataset presents significant variations in nuclear morphology, posing substantial challenges for training.
+Our experimental results, as shown in Table II (b), demonstrate that DAWN outperforms the leading method, WeakSAM, with improvements of 5.0% and 2.8% in DICE and AJI metrics, respectively.
+Additionally, DAWN shows notable gains over the second-best approach, achieving improvements of approximately 10%, 6%, and 8% in DICE, AJI, and PQ metrics, respectively.
 We perform visual comparison experiments, and the results are shown in the third and fourth rows of Fig.
 5.
-The results of WeakSeg, MaskGANet, C2FNet, and PENet show that these methods exhibit poorer segmentation results when faced with significant morphological variations of nuclei.
+The segmentation results of WeakSeg, MaskGANet, and C2FNet show that these methods exhibit poorer segmentation results when faced with significant morphological variations of the nuclei.
 In contrast, our method still achieves outstanding performance.
 We attribute this phenomenon to the fact that two-stage methods struggle to adapt their pseudo-labels to the variations in staining and morphology, leading to lower segmentation performance.
-However, our method utilizes a domain transfer strategy that leverages prior knowledge from the source domain and introduces a progressive domain adaptation method to adapt to the target domain’s image characteristics, resulting in better segmentation results.
+However, our method utilizes a domain transfer strategy that leverages prior knowledge from the source domain and introduces a progressive domain adaptation method to adapt to the target domain’s image characteristics.
 IV-C3 Comparisons on Colorectal Cancer Lizard and ConSeP
-We further validate the performance of our proposed method on colorectal cancer pathological images.
-From the Lizard results in Table II (c), we observe that existing weakly supervised methods perform poorly, especially C2FNet, MaskGANet, and PseEdgeNet, with PQ scores of only 0.134, 0.163, and 0.105, respectively.
-The phenomenon is because the Lizard dataset is sampled at the magnification of 20, resulting in densely distributed and small-sized nuclei, posing challenges for the segmentation task.
-Nevertheless, our method still achieves improvements compared to the best-performing WeakSeg.
-We observe improvements of 1.4% and 2.2% for DICE and AJI metrics, respectively.
-Additionally, from the fifth and sixth rows of Fig.
-5, it can be seen that the nuclei in the Lizard dataset are small and exhibit dense distribution, and most models struggle to produce satisfactory results.
-However, our approach can better identify the nuclei compared to other methods.
-From the ConSeP results in Table II (d), it can be observed that our method achieves significant improvements compared to other weakly supervised models.
-Compared to C2FNet, our DAWN model achieves improvements of 5.6% and 4.5% in terms of DICE and AJI, respectively.
-From the visual results in the last two rows of Fig.
-5, it can be seen that the nuclei in the ConSeP dataset often have blurry morphology and dense distribution.
-Other methods struggle to identify the boundaries and distinguish individual nuclei instances accurately.
-In contrast, our method partially addresses this issue and somewhat improves the segmentation performance.
+We further validate the performance of our proposed method in pathological images of colorectal cancer.
+As shown by the results of the Lizard dataset in Table II (c), existing weakly supervised methods, particularly C2FNet, MaskGANet, and PENet, perform poorly, with PQ scores of 0.134, 0.163, and 0.105, respectively.
+This underperformance can be attributed to the fact that the Lizard dataset is sampled at a magnification of ×$\times 20$, resulting in densely distributed and small nuclei, which presents significant challenges for segmentation.
+Despite these difficulties, our method achieves improvements over the best-performing method, WeakSeg, with increases of 1.4% and 2.2% in DICE and AJI metrics, respectively.
+Furthermore, as illustrated in the fifth and sixth rows of Fig.
+5, our method maintains excellent segmentation performance even with smaller nuclei.
+In contrast, the segmentation results of DDTNet, which performed well on the TNBC and CryoNuSeg datasets, are unsatisfactory on Lizard.
+This highlights that our method is less affected by the sampling magnification compared to other models.
+As shown in the ConSeP results in Table II (d), our proposed DAWN model achieves significant improvements over other weakly supervised models.
+Specifically, DAWN outperforms C2FNet with gains of 5.6% in DICE and 4.5% in AJI.
+The visual results in the last two rows of Fig.
+5 highlight that the nuclei in the ConSeP dataset often exhibit blurry morphology and are densely distributed, making it difficult for other methods to accurately delineate boundaries and distinguish individual nuclei instances.
+In contrast, our method addresses these challenges more effectively, leading to improved segmentation performance.
 Figure 5: The visualization comparison between weakly supervised methods and our proposed DAWN.
 IV-D Comparison with Fully Supervised Methods
-To compare the performance difference between our proposed weakly supervised segmentation method and existing fully supervised segmentation methods, we compare our DAWN with six state-of-the-art models, including U-Net [8], HoverNet [30], CDNet [32], DCAN [33], Micro-Net [68], and Dist [61].
+To compare the performance difference between our proposed weakly supervised segmentation method and existing fully supervised segmentation methods, we compare our DAWN with seven state-of-the-art models, including U-Net [10], HoverNet [33], CDNet [35], DCAN [32], Micro-Net [75], Dist [68], and AOST[76].
 In the experiment, we use PanNuke as the source domain and TNBC and ConSeP as the target domains.
@@
 The results are shown in Table II.
-In the table, the best fully supervised method is bold with red text, and the second best fully supervised method is underlined with red text.
+In the table, the best fully supervised method is bold, and the second best fully supervised method is underlined.
 IV-D1 Quantitative Comparisons
-From Table II, it can be seen that our method achieves comparable performance to fully supervised methods on the TNBC dataset.
-In the DICE, our method outperforms the best-performing CDNet by 0.4% and significantly surpasses other traditional supervised methods such as U-Net and DCAN.
-On the CryoNuSeg and ConSeP datasets, although our method does not achieve the best performance, it is only 1.8% and 0.4% lower than the second-best HoverNet method in terms of AJI.
-However, our method still performs poorly on the Lizard dataset.
-While it outperforms U-Net and DCAN, there is still a performance gap compared to other methods.
-We attribute this phenomenon to the limitations of our domain adaptation method when dealing with densely distributed nuclei data, and using point annotations for domain transfer still poses challenges.
+As shown in Table II, our method achieves performance comparable to fully supervised methods on the TNBC dataset.
+Specifically, our method surpasses the best-performing CDNet by 0.9% in DICE and significantly outperforms other traditional supervised methods such as U-Net and DCAN.
+On the CryoNuSeg and ConSeP datasets, although our method does not achieve the top performance, it is only 1.8% and 0.4% lower than the second-best HoverNet in terms of AJI.
+However, our method underperforms on the Lizard dataset.
+While it outperforms U-Net and DCAN, a performance gap remains compared to other methods.
+We attribute this to the inherent limitations of weakly supervised tasks, as using point annotations for domain transfer continues to present challenges.
 IV-D2 Qualitative Comparisons
@@
 6.
+In the figure, blue stands for correctly true positive pixels, red stands for false positive pixels and green stands for false negative pixels.
 From the figure we can see our weakly supervised segmentation method achieves comparable performance with supervised methods.
@@
 Blue stands for correctly true positive pixels, red stands for false positive pixels and green stands for false negative pixels.
-IV-E Domain Adaptation Evaluation
-We conduct a domain adaptation comparison with three domain adaptation methods: SIFA[17], PDAM[18] and SFDA-CellSeg[69]. The comparison experiment is designed from three perspectives.
-IV-E1 PanNuke as Source dataset
-To compare the different domain adaptation methods, we use a large pan-cancer dataset, PanNuke, as the source domain data and TNBC (breast cancer) and CryoNuSeg (pan-cancer) as the target domain data for the experiment.
-The experiment results are shown in Table III.
-From this table, we can observe that our method achieves significant improvements both on TNBC and CryoNuSeg datasets.
-Regarding the AJI metric, our method outperforms the second-best method by 4.63% and 4.49% on TNBC and CryoNuSeg, respectively.
-TABLE III: The domain adaptation comparison experiments from source domain PanNuke to target domain TNBC and CryoNuSeg.
+TABLE III: The domain adaptation comparison experiments from PanNuke dataset to TNBC and CryoNuSeg datasets.
 […]
@@
 SFDA-CellSeg [77] 0.8340 0.6817 0.8025 0.7933 0.6715
+ADPL[53] 0.8409 0.7122 0.8193 0.8117 0.6746
+CyCADA[54] 0.8093 0.6712 0.8249 0.7677 0.6343
+DDMRL[55] 0.8141 0.6869 0.8225 0.7763 0.6509
 Ours 0.8561 0.7280 0.8425 0.8202 0.6925
@@
 SFDA-CellSeg [77] 0.7912 0.4630 0.6121 0.7277 0.4480
+ADPL[53] 0.7847 0.4801 0.6104 0.7333 0.4542
+CyCADA[54] 0.7993 0.4794 0.6208 0.7278 0.4534
+DDMRL[55] 0.7764 0.4692 0.5998 0.7194 0.4449
 Ours 0.8041 0.5079 0.6374 0.7439 0.4760
-TABLE IV: The domain adaptation comparison experiments from source domain MoNuSeg to target domain TNBC and CryoNuSeg.
+IV-E Domain Adaptation Evaluation
+We conduct a domain adaptation comparison with six domain adaptation methods: SIFA[19], PDAM[20] and SFDA-CellSeg[77], ADPL[53], CyCADA[54], and DDMRL[55].
+The domain adaptation experiments are designed from the following three perspectives.
+IV-E1 PanNuke as Source dataset
+To compare the different domain adaptation methods, we use a large pan-cancer dataset, PanNuke, as the source domain data and TNBC (breast cancer) and CryoNuSeg (pan-cancer) as the target domain data for the experiment.
+The experiment results are shown in Table III.
+From this table, we can observe that our method achieves significant improvements both on TNBC and CryoNuSeg datasets.
+Regarding the AJI metric, our method outperforms the second-best method by 1.58% and 2.78% on TNBC and CryoNuSeg, respectively.
+TABLE IV: The domain adaptation comparison experiments from MoNuSeg dataset to TNBC and CryoNuSeg datasets.
 […]
@@
 SFDA-CellSeg [77] 0.8007 0.6597 0.7404 0.7484 0.6301
+ADPL[53] 0.8169 0.6723 0.7677 0.7648 0.6501
+CyCADA[54] 0.8204 0.6619 0.7549 0.7627 0.6480
+DDMRL[55] 0.8211 0.6544 0.7577 0.7612 0.6419
 Ours 0.8294 0.6933 0.8074 0.8085 0.6679
@@
 SFDA-CellSeg [77] 0.7687 0.4521 0.6004 0.7133 0.4287
+ADPL[53] 0.7527 0.4644 0.5998 0.7107 0.4250
+CyCADA[54] 0.7622 0.4548 0.6010 0.7109 0.4257
+DDMRL[55] 0.7601 0.4547 0.5989 0.7034 0.4241
 Ours 0.7864 0.4845 0.6122 0.7216 0.4498
@@
 The experiment results are shown in Table IV.
-When using MoNuSeg as the source domain data, the AJI metric of our method surpasses the second-best method by 3.36% and 3.24% on TNBC and CryoNuSeg target data, respectively.
+When using MoNuSeg as the source domain data, the AJI metric of our method surpasses the second-best method by 2.10% and 1.91% on TNBC and CryoNuSeg target data, respectively.
 This indicates that our method is effective on small-scale datasets.
 Besides, comparing the results in Table III and Table IV, the latter’s performance significantly decreases.
-On the CryoNuSeg data, the DICE metric for the three comparison methods decreased by 0.33%, 2.06%, and 2.25%, while our proposed method decreased by only 1.77%, showing a smaller decline compared to the PDAM and SFDA-CellSeg.
+On the CryoNuSeg data, the DICE of the ADPL and CyCADA decreased by 3.20% and 3.71%, while our proposed method decreased by only 1.77%, showing a smaller decline compared to the ADPL and CyCADA.
 The phenomenon suggests that our method is less affected by the data scale disturbance.
-IV-E3 Cross-cancer Domain Adaptation
-We conduct cross-cancer domain adaptation experiments on the ConSeP (colon cancer) and TNBC (breast cancer) datasets, and the experiment results are shown in Table V.
-We can observe that our method achieves the best performance among these methods.
-Particularly on the TNBC, the AJI surpasses the second-best method by 4.20%.
-These experiments demonstrate the significant advantage of our method.
-Based on this phenomenon, we can conclude that our method is least affected by the source domain data.
-Our method can still exploit essential prior information for different cancer types to enhance the segmentation performance.
-Figure 7: The comparison of before and after domain adaptation.
-Blue circles indicate instances which are not recognition in source only method; Red and white circles represent accurate instance segmentation in our DAWN; Yellow arrow represents more smoother boundary of our DAWN, which is enlarged in bottom-right corner.
-TABLE V: The cross-cancer domain adaptation experiments from source domain ConSeP (TNBC) to target domain TNBC (ConSeP).
+Figure 7: The visualization comparisons between three domain adaptation methods and our proposed DAWN from PanNuke dataset to TNBC dataset.
+TABLE V: The cross-cancer domain adaptation experiments from ConSeP (TNBC) dataset to TNBC (ConSeP) dataset.
 […]
@@
 SFDA-CellSeg [77] 0.7582 0.6081 0.6033 0.7322 0.5921
+ADPL[53] 0.7696 0.6231 0.6547 0.7597 0.6120
+CyCADA[54] 0.7654 0.6209 0.6383 0.7404 0.6001
+DDMRL[55] 0.7585 0.6199 0.6217 0.7383 0.5997
 Ours 0.8285 0.6634 0.7894 0.7855 0.6453
@@
 SFDA-CellSeg [77] 0.6913 0.4217 0.4484 0.6520 0.3817
+ADPL[53] 0.7174 0.4437 0.4824 0.6634 0.3883
+CyCADA[54] 0.7274 0.4430 0.4746 0.6565 0.3943
+DDMRL[55] 0.7107 0.4349 0.4680 0.6443 0.3814
 Ours 0.7743 0.4607 0.5549 0.6981 0.4168
+IV-E3 Cross-cancer Domain Adaptation
+We conduct cross-cancer domain adaptation experiments on the ConSeP (colon cancer) and TNBC (breast cancer) datasets, and the experiment results are shown in Table V.
+We can observe that our method achieves the best performance among these methods.
+Particularly on the TNBC, the AJI surpasses the second-best method by 4.03%.
+These experiments demonstrate the significant advantage of our method.
+Based on this phenomenon, we can conclude that our method is least affected by the source domain data.
+Our method can still exploit essential prior information for different cancer types to enhance the segmentation performance.
 IV-E4 Qualitative Comparisons
-We provided visualization comparisons of different methods in Fig.
-7 and Fig.
+We provide visualization comparisons of segmentation results between domain adaptation methods in Fig.
+7.
+First, our model exhibits better nuclei segmentation capability compared to the source domain method, especially for missed and overlapping nuclei.
+In contrast, SIFA[19] and PDAM[20] exhibit poor segmentation performance on some overlapping nuclei.
+These results further demonstrate the effectiveness of our proposed method in domain adaptation, whether from pan-cancer to specific cancer types or from pan-cancer to other pan-cancer datasets.
+IV-E5 Detection performance
+To evaluate the influence of our proposed auxiliary detection branch on nuclei detection, we conduct a domain adaptation comparison experiment on four target datasets in three metrics, i.e., Recall, Precision, and F1.
+The experimental results are shown in Table VI.
+As seen from the table, when using the detection branch to assist domain transfer, the Recall has been significantly improved without reducing F1, which means that the auxiliary task can help the segmentation network to identify the nucleus more completely.
+This conclusion can also be verified from the visualization results in Fig.
 8.
-From Fig.
-7, comparing with only using source domain method, our domain adaptation method exhibits improved recognition capability for overlapping and missed nuclei by incorporating an output of detection network in CPL.
-In Fig.
-8, our results are closest to the nuclei segmentation ground truth.
-In contrast, SIFA[17] and PDAM[18] exhibit poorer segmentation performance on some overlapping nuclei.
-These results further demonstrate the effectiveness of our proposed method in domain adaptation, whether from pan-cancer to specific cancer types or from pan-cancer to other pan-cancer datasets.
-Figure 8: The visualization comparisons between three domain adaptation methods and our proposed DAWN from PanNuke dataset to TNBC dataset.
-TABLE VI: The ablation experiments of CFC and IST modules on TNBC, CryoNuSeg, Lizard and ConSeP datasets.
+The figure’s red $\star$ represents the actual nuclei location.
+Orange $\bullet$” and blue “+” represent the predicted location before and after domain adaptation.
+The figure shows that after adding the detection task, the predicted results are highly coincident with the actual labels, indicating that our detection task helps enhance the nuclei recognition ability of the segmentation task.
+TABLE VI: The detection results before and after domain adaptation.
+Datasets Stage Recall Precision F1
+TNBC Bef. 0.7647 0.8949 0.8205
+Aft. 0.8605 (0.10$\uparrow$) 0.7870 (0.11$\downarrow$) 0.8221 (0.002$\uparrow$)
+CryoNuSeg Bef. 0.4637 0.7522 0.5737
+Aft. 0.6635 (0.20$\uparrow$) 0.7326 (0.02$\downarrow$) 0.6963 (0.12$\uparrow$)
+Lizard Bef. 0.7804 0.8373 0.7958
+Aft. 0.8029 (0.02$\uparrow$) 0.8109 (0.03$\downarrow$) 0.8069 (0.01$\uparrow$)
+ConSeP Bef. 0.5563 0.7893 0.6526
+Aft. 0.7878 (0.23$\uparrow$) 0.5999 (0.19$\downarrow$) 0.6812 (0.03$\uparrow$)
+Figure 8: The nuclei detection results. (a) represents the detection results of source only method; (b) represents the detection results of our DAWN.
+TABLE VII: The ablation experiments of CFC and IST modules on TNBC, CryoNuSeg, Lizard and ConSeP datasets.
 Data CFC IST DICE AJI DQ SQ PQ
@@
 IV-F Ablation study
-IV-F1 Evaluation of CFC and IST
-To prove the effectiveness of CFC and IST components, we perform ablation studies with results listed in Table VI.
+IV-F1 Evaluation of module designs
+To prove the effectiveness of CFC and IST components, we perform ablation studies with results listed in Table VII.
 On the TNBC and Lizard datasets, the five metrics have noticeable improvements whether CFC or IST is added.
@@
 Our ablation study proves that the CFC and IST promote domain adaptation ability and enhance nuclei segmentation performance on weakly supervised scenarios.
-TABLE VII: The mask fusion and distance filter ablation experiments on TNBC dataset.
+To prove that mask fusion and distance filter effectively improve segmentation performance, we perform an ablation study and parameter-sensitive experiments on the TNBC dataset.
+The results are shown in Table VIII.
+From the result, we can see that, whether removing mask fusion or distance filter, the performance decreases evidently, which indicates the two components are effective in the CPL module because they can enhance the quality of pseudo-labels.
+Figure 9: The iteration ablation study on four datasets. The x-axis represents the number of iterations, and the y-axis represents model performance.
+TABLE VIII: The mask fusion and distance filter ablation experiments on TNBC dataset.
 Methods PanNuke $\rightarrow$ TNBC
@@
 w/o distance 0.8357 0.7045 0.8258 0.8232 0.6814
-ours(d=10) 0.6541 0.4917 0.5771 0.7054 0.4170
-ours(d=25) 0.8561 0.7280 0.8425 0.8202 0.6925
-ours(d=40) 0.8332 0.7120 0.8313 0.8220 0.6850
-Figure 9: The iteration ablation study on four datasets. The x-axis represents the number of iterations, and the y-axis represents model performance.
-Figure 10: The iteration training loss on TNBC dataset. The x-axis represents the training step and the y-axis represent the training loss and F1 value.
-TABLE VIII: The detection results before and after domain adaptation.
-Datasets Stage Recall Precision F1
-TNBC Bef. 0.7647 0.8949 0.8205
-Aft. 0.8605 (0.10$\uparrow$) 0.7870 (0.11$\downarrow$) 0.8221 (0.002$\uparrow$)
-CryoNuSeg Bef. 0.4637 0.7522 0.5737
-Aft. 0.6635 (0.20$\uparrow$) 0.7326 (0.02$\downarrow$) 0.6963 (0.12$\uparrow$)
-Lizard Bef. 0.7804 0.8373 0.7958
-Aft. 0.8029 (0.02$\uparrow$) 0.8109 (0.03$\downarrow$) 0.8069 (0.01$\uparrow$)
-ConSeP Bef. 0.5563 0.7893 0.6526
-Aft. 0.7878 (0.23$\uparrow$) 0.5999 (0.19$\downarrow$) 0.6812 (0.03$\uparrow$)
-Figure 11: The nuclei detection results. (a) represents the detection results of source only method; (b) represents the detection results of our DAWN.
-IV-F2 Evaluation of Hyper-parameter in CPL
-To prove that mask fusion and distance filter effectively improve segmentation performance, we perform an ablation study and parameter-sensitive experiments on the TNBC dataset.
-The results are shown in Table VII.
-From the result, we can see that, whether removing mask fusion or distance filter, the performance decreases evidently, which indicates the two components are effective in the CPL module because they can enhance the quality of pseudo-labels.
-Besides, we can see that the model is sensitive to the distance filter method.
-When the distance is small, the performance decreases because it destroys the nuclear shape.
-When the distance is enlarged, the performance is close to the result without a distance filter.
-Overall, the mask fusion and distance filter in CPL have also been proved to improve performance.
+ours 0.8561 0.7280 0.8425 0.8202 0.6925
+TABLE IX: The hyper-parameter $r_{1}$ ablation experiments on TNBC dataset.
+Methods PanNuke $\rightarrow$ TNBC
+DICE AJI DQ SQ PQ
+$r_{1}$=5 0.8147 0.6828 0.7914 0.7784 0.6554
+$r_{1}$=11 0.8561 0.7280 0.8425 0.8202 0.6925
+$r_{1}$=20 0.8323 0.7024 0.8191 0.7923 0.7022
+IV-F2 Evaluation of Hyper-parameter
+First, by observing Table IX, there is a significant decrease in the performance of the model when the radius range of the nuclei is altered.
+Whether $r_{1}$ is set to $5$ or $20$, the Dice coefficient drops by more than 2 %.
+This indicates that the setting of nucleus radius has a substantial impact on the model and that configuring parameters based on the radius from the validation set is advantageous.
+Secondly, observing Tables X and XI it can be seen that there is no significant variation in model performance when modifying the standard deviation $\sigma$ and $\theta$ of the nuclear radius.
+In fact, in some metrics, the model even outperforms the reported results in the main text; for example, when $\sigma$ = 4, the SQ metric improves by 0.2%.
+However, this variation is not substantial, indicating that the model parameters are less affected by $\sigma$ and $\theta$.
+Lastly, we also include additional sensitivity analyses of distance $d$ as presented in Table XII.
 IV-F3 Evaluation of Iterative Training
@@
 9.
-In this figure, the x-axis represents the number of iterations, corresponding to the parameter ”r” in Algorithm 1, while the y-axis represents the values of the DICE, AJI, and PQ metrics.
+In this figure, the x-axis represents the number of iterations, corresponding to the parameter ”I” in Algorithm 1, while the y-axis represents the values of the DICE, AJI, and PQ metrics.
 The results show that on the TNBC and Lizard datasets, the model’s performance shows an increasing trend as the number of training iterations increases.
-Specifically, the TNBC dataset achieves optimal performance at r=3, while the Lizard dataset performs best at r=4.
+Specifically, the TNBC dataset achieves optimal performance at I=3, while the Lizard dataset performs best at I=4.
 In contrast, on the CryoNuSeg and ConSeP datasets, the model’s performance remains relatively stable after the second round of training.
-Consequently, we select the optimal performance according to r as the segmentation result.
+Consequently, we select the optimal performance according to I as the segmentation result.
 Furthermore, Fig.
-10 presents the loss function and F1 curve for three iterations of training on the TNBC dataset, with each iteration consisting of 20 training epochs.
+10 presents the loss function and the F1 curve for three training iterations in the TNBC dataset, each iteration consisting of 20 training epochs.
 From this figure, it can be observed that as training progresses, the model’s loss steadily decreases.
 Correspondingly, the F1 score increases with training and performs best in the final iteration.
-IV-F4 Evaluation of Domain Adaptation
-To evaluate the influence of our proposed auxiliary detection branch on nuclei detection ability, we conduct a domain adaptation comparison experiment on four target datasets in three metrics, i.e., Recall, Precision, and F1.
-The experimental results are shown in Table VIII.
-As seen from the table, when using the detection branch to assist domain transfer, the Recall has been significantly improved without reducing F1, which means that the auxiliary task can help the segmentation network to identify the nucleus more completely.
-This conclusion can also be verified from the visualization results in Fig.
+TABLE X: The hyper-parameter $\sigma$ ablation experiments on TNBC dataset.
+Methods PanNuke $\rightarrow$ TNBC
+DICE AJI DQ SQ PQ
+$\sigma$=2.0 0.8447 0.7120 0.8394 0.8191 0.7022
+$\sigma$=2.75 0.8561 0.7280 0.8425 0.8202 0.6925
+$\sigma$=4.0 0.8402 0.7221 0.8334 0.8227 0.6896
+TABLE XI: The hyper-parameter $\theta$ ablation experiments on TNBC dataset.
+Methods PanNuke $\rightarrow$ TNBC
+DICE AJI DQ SQ PQ
+$\theta$=0.1 0.8489 0.7233 0.8510 0.8212 0.6917
+$\theta$=0.2 0.8561 0.7280 0.8425 0.8202 0.6925
+$\theta$=0.3 0.8550 0.7226 0.8390 0.8226 0.6898
+TABLE XII: The hyper-parameter $d$ ablation experiments on TNBC dataset.
+Methods PanNuke $\rightarrow$ TNBC
+DICE AJI DQ SQ PQ
+$d$=10 0.6541 0.4917 0.5771 0.7054 0.4170
+$d$=25 0.8561 0.7280 0.8425 0.8202 0.6925
+$d$=40 0.8332 0.7120 0.8313 0.8220 0.6850
+Figure 10: The iteration training loss on TNBC dataset. The x-axis represents the training step and the y-axis represent the training loss and F1 value.
+Figure 11: The limitation of our proposed DAWN when facing nuclei with significant morphological variations.
+V Conclusion and Limitation
+In this paper, to reduce the reliance on manual annotations for nuclei segmentation, we propose a domain-adaptive weakly supervised segmentation method, termed DAWN.
+Our approach enables the transfer of a model trained on public data (source domain) to arbitrary external datasets (target domain) without requiring fine pixel-level annotations.
+This domain adaptation is guided by an auxiliary detection task.
+DAWN employs a Consistent Feature Constraint (CFC) module to improve domain transfer efficiency at the feature level.
+Additionally, we introduce a combined pseudo-label optimization method (CPL) and an interactive supervision training (IST) strategy to enhance domain adaptation at the task level.
+Ablation experiments confirm the effectiveness of these components in improving nuclei segmentation performance.
+Despite these advancements, DAWN exhibits limitations when handling nuclei with significant intra-image morphological variations, such as large spindle-shaped nuclei coexisting with small, circular ones, as illustrated in Fig.
 11.
-The figure’s red $\star$ represents the actual nuclei location.
-Orange $\bullet$” and blue “+” represent the predicted location before and after domain adaptation.
-The figure shows that after adding the detection task, the predicted results are highly coincident with the actual labels, indicating that our detection task helps enhance the nuclei recognition ability of the segmentation task.
-V Conclusion
-To reduce the dependence of nuclei segmentation on manual annotation, we propose a domain-adaptive weakly supervised segmentation method named DAWN.
-Our method can transfer the model trained on the public data (source domain) to arbitrary external data (target domain) without fine pixel-level annotation.
-For this purpose, the domain adaptation of the segmentation network is guided by an auxiliary detection task.
-In DAWN, a consistent feature constraint (CFC) module is used to enhance the domain transfer efficiency at the feature level.
-In addition, a combined pseudo-labels optimization (CPL) method and interactive supervision training (IST) method are proposed to enhance the domain adaptation ability at the task level.
-The ablation experiments prove these components are effective in boosting the nuclei segmentation performance.
-Furthermore, our comparison experiments show that our DAWN outperforms existing weakly supervised methods and achieves similar performance compared with supervised methods.
+This leads to occasional segmentation errors, particularly under-segmentation of nuclei with atypical shapes.
+To address these challenges, future work will focus on developing enhanced interaction mechanisms between the segmentation and detection networks.
+Specifically, we aim to introduce a multi-task learning framework that simultaneously optimizes segmentation and shape detection, allowing the model to better differentiate and capture varying nuclear shapes by leveraging contextual and spatial information across scales.
+Additionally, we plan to incorporate a region-aware feature aggregation module, which will allow the model to adaptively focus on regions with high morphological diversity.
+This approach aims to improve segmentation accuracy for nuclei with atypical geometries by explicitly modeling the relationships between various shapes within the same image.
+By combining these stratesgies, we aim to improve DAWN’s ability to segment diverse and complex nuclei, pushing the boundaries of domain adaptation in histopathology image analysis.